### v2

# Global-Local Interface with Selective Direct and Singularity-Avoiding Motion Mapping for Intuitive Teleoperation

Abstract
This paper presents the Global–Local (G–L) Teleoperation Interface, a hierarchical framework that enhances human–robot interaction by decoupling large-scale manipulator positioning from fine end-effector manipulation. The global component enables efficient workspace traversal, while the local component facilitates precise and dexterous maneuvering. To address kinematic singularities on the slave side—particularly during local fine manipulation—we propose a singularity-avoiding motion mapping strategy that improves both stability and intuitiveness. Additionally, we introduce the concept of an operational Jacobian to formally analyze the smoothness of slave joint motions under local control. The G–L interface is realized in two variants: Direct Mapping and Singularity-Avoiding Mapping. Both implementations are validated through hardware experiments involving tasks that demand precision, extended linear motion, and large orientation changes. Results demonstrate significant improvements in task success rate, execution efficiency, and user experience compared to conventional systems relying solely on global or local control. These findings indicate that the G–L interface offers a robust and scalable foundation for intuitive and versatile teleoperation across complex tasks and unstructured environments.
Index Terms:
Teleoperation, Human-Robot Interface, Kinematic Mapping, Singularity Avoidance
Figure 1: Comparison of teleoperation interfaces across different system configurations. Our proposed G-L interface (both bimanual and single-arm variants) offers broader joint-space coverage and higher precision, while retaining compatibility with complex end-effectors.
I Introduction
Teleoperation serves as a critical human–robot interface paradigm, enabling remote robotic operation in scenarios where human manual manipulation is insufficient or unsafe, such as surgery, disaster response, and exploration in hazardous or unstructured environments [1, 2, 3, 4, 5, 6, 7, 8]. It also plays an increasingly essential role in data-driven robotic learning, where high-quality human demonstration datas are crucial for imitation learning and multimodal policy training [9, 10, 11, 12, 13, 14, 15, 16].
Despite extensive research, existing teleoperation frameworks exhibit fundamental trade-offs among dexterity, workspace range, and intuitiveness. One common approach employs six-DoF tracking of the operator’s hand using floating controllers or vision-based systems [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. These enable freehand dexterity but often lack physical correspondence with the slave robot’s workspace, leading to non-intuitive motion mapping and vulnerability against manipulator singularities.
Alternatively, kinematically matched master devices replicate the structure of the slave robot using a scaled joint-link chain, or in some cases, by employing an identical slave device as the master itself [27, 28, 29, 30, 31]. This improves mapping of reachable configurations and collision awareness. However, such systems often constrain the operator’s movements and suffer from encoder detent torque, mechanical friction, and limited fidelity during fine manipulation.
To overcome these limitations, we propose the Global–Local (G–L) Teleoperation Interface (Fig. 1), a framework that hierarchically decouples human input into two synergistic control pathways: a global controller for large-scale positioning and a local controller for high-precision end-effector manipulation. This modular design fuses joint space awareness and dexterous control while minimizing the drawbacks of traditional systems. The G–L interface is a general teleoperation framework that supports flexible combinations of global and local master devices, allowing system configurations to be tailored to diverse tasks and hardware platforms. In addition, to facilitate more intuitive orientation control and avoid the impact of manipulator singularity while the local component is activated, we introduce a practical singularity-avoiding motion mapping strategy targeting the Pieper kinematics structure [32] between the local component motion and slave system’s joints. A concept termed operational Jacobian is proposed to give a rigorous mathematical argument of why this singularity-avoiding motion mapping ensures smooth slave joint motions at manipulator singularities. This concept can also be generalized to analyze the teleoperation control on slave systems other than Pieper structures.
Experimental validation and demonstration of the G-L interface utilized two different implementations defined by the mapping from the local component to the slave system’s joints. The first, designated Direct Mapping, maps the scaled Cartesian motion of the local component directly to the slave end-effector. This preserves six degrees of freedom for local manipulation but is characterized by less intuitive orientation mapping and the absence of a singularity avoidance feature. In contrast, the second implementation, Singularity-Avoiding Mapping, employs the aforementioned manipulator singularity avoidance technique. Although this method reduces the available degrees of freedom by one, it provides more intuitive orientation mapping and robustly avoids singularities within its designated workspace.
The contributions of this work are summarized as follows.
  1. 1.
    Proposal of the G-L Teleoperation Interface: A hierarchical architecture that decouples large-scale positioning from fine end-effector manipulation, designed to enhance operational flexibility and task-specific performance.
  2. 2.
    Implementation of Two Interface Realizations: Development and demonstration of the G-L concept through two distinct hardware implementations: the Direct Mapping Realization and the Singularity-Avoiding Realization. The later introduces a specialized motion mapping strategy for the local control component, aimed at improving orientation control intuitiveness and ensuring stable operation near slave manipulator singularities.
  3. 3.
    Introduction of the Operational Jacobian: Proposal of the operational Jacobian as a mathematical tool to rigorously analyze the smoothness of the slave system’s joint motion during teleoperation.
  4. 4.
    Experimental Validation and Comparative Analysis: Comprehensive validation of the G-L interface through experiments involving precision manipulation, large linear movements, and significant angular adjustments. Results demonstrate quantifiable improvements in success rates, operational efficiency, and user experience compared to conventional single-mode (global-only or local-only) systems.
The remainder of this paper is organized as follows. Section II introduces the proposed Global-Local (G-L) teleoperation interface and its key design criteria. Section III details the direct mapping realization, while Section IV presents the singularity-avoiding mapping strategy. Section V describes the operational Jacobian tool. Experimental validations across various manipulation tasks are discussed in Section VI. Finally, Section VII concludes the paper with a summary of contributions, limitations, and directions for future work.
II Global-Local Teleoperation Interface
To be considered a valid implementation of the G-L Interface, a decoupled master system must adhere to the following criteria. First, the global component must provide an intuitive and comprehensive representation of the slave system’s collision configuration. It should enable rapid, large-amplitude adjustments of the slave’s pose, prioritizing efficient workspace navigation over fine manipulation or dexterity. Second, the local component must facilitate precise, fine-grained control of the slave’s end-effector, enabling the operator to execute tasks requiring dexterity and accuracy. It should prioritize the seamless transfer of the operator’s manipulation skills to the slave system, with less emphasis on reflecting the slave’s overall configuration or controlling all DoFs. In the remaining of this paper, without further clarification, our discussions are limited to slave systems with serial connected structure.
Employing a scaled replica of the slave system, a common strategy in teleoperation [9, 30], effectively meets the requirements for the global component of our system. However, this implementation differs from existing scaled replica teleoperation systems in two ways. Firstly, the global component does not require exact kinematic correspondence with the slave system’s revolute axes, as its primary function is to enable rapid and coarse motions. It only needs to provide an approximation of the current slave system’s collision configuration, allowing the human operator to avoid potential collisions during coarse adjustments of the slave’s pose. Secondly, the global component need not encompass all Degrees of Freedom (DoFs) of the slave system. It is generally observed that DoFs closer to the base exert a more substantial influence on both the robot’s end-effector position and its overall collision state compared to DoFs nearer the end-effector. Consequently, it may be sufficient for the global component to replicate only these more influential, base-proximal DoFs.
Tracking the operator’s hand motion represents a suitable approach for implementing the local component of the interface. A key challenge in Global-Local (G-L) interface design is establishing an intuitive motion mapping between this local component and the slave system. Focusing on the widely applied Pieper kinematics manipulator structure, we propose two distinct motion mapping methods. The first, more direct method involves mapping the local component’s motion to the slave system’s end-effector motion and solving for the required joint poses using inverse kinematics. Owing to its straightforward implementation and preservation of all 6 DoFs for the slave end-effector, this technique is commonly adopted in robot data collection applications [25]. However, if the initial orientation of the local component does not align well with the slave end-effector’s starting orientation, this method can prove counter-intuitive for orientation control. Furthermore, manipulator singularities can adversely affect the usability of the local component, an issue detailed further in Section IV. The second method maps the three rotational DoFs of the local component to three mutually perpendicular slave joints and locks one slave DoF to mitigate certain rotational singularities. While this approach reduces the effective mapped DoFs by one, it provides considerably more intuitive orientation control.
Using the two local motion mapping approaches and with different number of DoFs the global components cover, we present the Direct Mapping and Singularity-Avoiding Mapping realizations of the G-L interface, which will be introduced in detail in sections III and IV. Both realizations share a challenge during the transition from the local to global component. A mismatch can arise between the pose of the scaled replica (global component) and the actual slave system configuration, potentially causing an abrupt motion in the slave upon returning to global control. To mitigate this, the global component assumes a ”slave” role during local component activation, mirroring the joint motions of the true slave system. This ensures pose synchronization between the global component and the slave, facilitating a smooth transition when switching back to global controls.
Fig. 2 illustrates a representative configuration of the proposed Global–Local (G–L) interface. Haptic devices are used to capture the operator’s fine hand motions, while a scaled-down 3-DOF joint-link chain captures the large range arm movement. Two pairs of foot pedals allow the operator to switch between global and local control modes.
Figure 2: A schematic diagram of G-L interface. The global component is a scaled replica of the slave robot and the local component is a haptic device. The two step pedals allows the operator to rapidly switch between the local and global component. If the local component is activated, the global component will temporary turn into a slave system that follows the actual slave system to avoid joint position mismatch when switching back to the global component.
III Direct Mapping of Local Component Motion
It is important to emphasize the key distinction between the Direct Mapping and Singularity-Avoiding Mapping strategies: they primarily differ in the local motion mapping method and the number of degrees of freedom allocated to the global component.
The global component of the Direct Mapping implementation covers the full six DoFs in the Pieper structure slave robot [33, 34, 35]. Activating the global component places the slave system in joint control mode, replicating the scaled replica’s joint positions within the limits of the motor capacities. Conversely, activating the local component switches the slave system to end-effector Cartesian control, where it tracks the motion of the haptic device.
To facilitate precise control of the end-effector, the slave system is designed to track only a portion of the local component motion. Denote the slave end-effector pose with $(\bm{p},\bm{R})$ where $\bm{p}\in\mathbb{R}^{3}$ is the position and $\bm{R}\in\mathbb{R}^{3\times 3}$ is the orientation represented by a rotational matrix. Let $\bm{q}$ be the joint poses of the slave robot, then we have $\bm{q}=\bm{F}(\bm{p},\bm{R})$ where $\bm{F}$ is the inverse kinematics of Pieper kinematics structure, which can be efficiently computed in closed form [32].
Every time the local component is activated, its starting pose $(\bm{p}^{m}_{0},\bm{R}^{m}_{0})$ as well as the slave end-effector’s starting pose $(\bm{p}^{s}_{0},\bm{R}^{s}_{0})$ will be recorded. Suppose at a certain moment $t$, the local component moves to pose $(\bm{p}^{m}_{t},\bm{R}^{m}_{t})$, we first compute the linear and rotational displacements since activation.
$\begin{split}\bm{dp}_{t}^{m}&=\bm{p}_{t}^{m}-\bm{p}_{0}^{m}\\ (\bm{l}^{m}_{t},\theta^{m}_{t})&=\bm{G}(\bm{dR}^{m}_{t})=\bm{G}(\bm{R}^{m,T}_{0}R^{m}_{t})\end{split}$ (1)
where the superscript $T$ stands for transpose and $\bm{G}$ maps the rotational matrix to axis-angle format. $\bm{l}_{t}$ is the normalized rotational axis direction and $\theta_{t}$ is the angle.
Given prespecified linear and rotational scaling factors $\alpha_{l},\alpha_{r}\in(0,1]$, the slave system will move its end-effector pose towards the scaled linear and rotational displacements of the haptic device. Mathematically, the target slave pose at moment $t$ is
$\begin{split}\bm{p}_{t}^{s}&=\bm{p}_{0}^{s}+\alpha_{l}\bm{dp}_{t}^{m}\\ \bm{R}_{t}^{s}&=\bm{R}_{0}^{s}\bm{G}^{-1}(\bm{l}^{m}_{t},\alpha_{r}\theta^{m}_{t})\end{split}$ (2)
with $\bm{G}^{-1}$ being the mapping that gets the rotational matrix from the axis-angle interpretation.
With the generated end-effector pose, the target joint poses of the slave robot at moment $t$ will be
$\bm{q}^{s}_{t}=\bm{F}(\bm{p}^{s}_{t},\bm{R}^{s}_{t})$ (3)
IV Singularity-Avoiding Mapping of Local Component Motion
Despite the preservation of full degrees of freedom of the end-effector, the aforementioned direct motion mapping has two major limitations.
  • It is particularly vulnerable against manipulator singularities, especially the wrist singularity in Pieper kinematics, which can happen everywhere inside the manipulator reachable workspace [36]. This demands human operators to pay extra attention on the slave joint poses other than the end-effector pose.
  • When the starting orientation of the local component and the slave end-effector have an obvious misalignment, controlling orientation using direct mapping can be overwhelming for human operators. This is because it is difficult for human operators to predict the slave end-effector’s orientation change given the motion of the local component. Additionally, they cannot intuitively discern which joints will be actuated and how much they will move.
This section focuses on the scenario when the slave system has a Pieper kinematic structure and proposes a practical local component motion mapping method to alleviate these limitations.
IV-A Pieper Kinematics and Translation-Rotation Decoupling
Pieper kinematics structure, also known as the 321 kinematics structure, is widely adopted in 6-joint industrial robot arms for its closed-form inverse kinematics. Fig. 3 is an illustration of the three most important characteristics of Pieper kinematics structure using AgileX Piper as an example.
  • The base joint (J1) and the first shoulder joint (J2) intersect and are perpendicular.
  • The two shoulder joints (J2 and J3) are parallel.
  • The three wrist joints (J4, J5 and J6) intersects at the same point $Q$, usually arranged in an intrinsic X-Y-X Euler Angles order.
Figure 3: (A) Pieper Kinematics Structure using Agile X Piper as an example. (B) $q_{5}=0^{\circ}$ wrist singularity, which is a typical manipulation pose and is also the singularity type we are focusing on avoiding.
It can be observed that the intersection point $Q$ is relatively close to the end-effector tip and its position is only affected by the poses of the base and shoulder joints. In addition, since the three wrist joints formulate an Euler angle combinations, they can dexterously adjust the end-effector orientation in a nonzero-measure set without actuating J1, J2 and J3, regardless of the Cartesian position of $Q$ [37]. With this observation, the following amendments are made to the Direct Mapping realization.
  • Translation of the local component now controls the linear motion of the intersection point $Q$, $\bm{p}_{t}^{Q}$, not the end-effector’s linear motion $\bm{p}_{t}^{s}$. Therefore, only J1, J2 and J3 will be actuated.
  • Rotation of the local component still controls the slave end-effector’s orientation $\bm{R}_{t}^{s}$, but only J4, J5 and J6 will be actuated.
Denote the starting pose of point $Q$ as $(\bm{p}_{0}^{Q},\bm{R}_{0}^{Q})$, the target position of point $Q$ at moment $t$ will be
$\bm{p}_{t}^{Q}=\bm{p}_{0}^{Q}+\alpha_{l}\bm{dp}_{t}^{m}$ (4)
The joint poses are now solved as
$\begin{split}\begin{bmatrix}q_{1,t}^{s},q_{2,t}^{s},q_{3,t}^{s}\end{bmatrix}^{T}&=\bm{F}_{123}(\bm{p}_{t}^{Q})\\ \begin{bmatrix}q_{4,t}^{s},q_{5,t}^{s},q_{6,t}^{s}\end{bmatrix}^{T}&=\bm{F}_{456}(\bm{R}_{t}^{Q,T}\bm{R}_{t}^{s})\end{split}$ (5)
where $\bm{F}_{123}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ is the mapping from the $Q$ point position to J1, J2 and J3 joint poses. $\bm{F}_{456}:\mathbb{R}^{3\times 3}\rightarrow\mathbb{R}^{3}$ maps from rotation matrix to J4, J5 and J6 joint poses. $\bm{R}_{t}^{Q}$ is the current orientation of point $Q$.
This decoupling process of the local component’s translation and rotation partially reduces the corresponding number of joints of each local component DoF to less than or equal to three. Although this increases control intuitiveness for human operators, it still suffers from manipulator singularities. In section IV-B, we discuss further amendments that can help avoid manipulator singularities.
Figure 4: Visualization of Singularity Avoidance Motion Mapping. The digital pen in the center indicates the direction where the local component is pointing to. (A) Row angle $d\theta_{t}^{m}$ maps to J1 motion. Notice that under motion mapping (11), the digital pen does not necessaily align with the X-axis at the beginning. (B) Pitch angle $d\phi_{t}^{m}$ maps to J5 motion. (C) Yaw angle $d\psi_{t}^{m}$ maps to J6 motion.
IV-B Singularities in Pieper Kinematics
Following notations in [36], the motion mapping from the configuration space to the operational space has the following form.
$\bm{v}^{e}_{t}=\bm{J}(\bm{q}_{t})\dot{\bm{q}_{t}}$ (6)
An important difference here is that $\bm{v}^{e}_{t}$ now consists of the linear velocity of point $Q$ and the angular velocity of the end-effector $\bm{v}^{e}_{t}=\begin{bmatrix}\dot{\bm{p}}^{Q}_{t}&\bm{\omega}^{e}_{t}\end{bmatrix}$. Since Pieper kinematics structure has six independent joints, the Jacobian can be naturally partitioned into $(3\times 3)$ blocks.
$\bm{J}(\bm{q}_{t})=\begin{bmatrix}\bm{J}_{11}(\bm{q}_{t})&\bm{J}_{12}(\bm{q}_{t})\\ \bm{J}_{21}(\bm{q}_{t})&\bm{J}_{22}(\bm{q}_{t})\end{bmatrix}$ (7)
Manipulator singularities are the joint configurations $\bm{q}$ where the matrix $\bm{J}(\bm{q})$ presents a zero singular value. When $\bm{J}(\bm{q})$ is a square matrix, this is equivalent to having a zero determinant, $\det{\bm{J}(q)}=0$. Since the position $\bm{p}_{t}^{Q}$ is independent of J4-J6, we know that $\bm{J}_{12}(\bm{q}_{t})=\bm{O}_{3}$. As a result, manipulator singularities can be identified from the following equation.
$\det\bm{J}_{11}(\bm{q})\det\bm{J}_{22}(\bm{q})=0$ (8)
Scenarios when $\det\bm{J}_{11}(\bm{q})=0$ is often referred to as arm singularities since it mainly affects the linear velocity and $\det\bm{J}_{22}(\bm{q})=0$ is often referred to as wrist singularities since it mainly affect angular velocities. $\bm{J}_{11}(\bm{q})$ and $\bm{J}_{22}(\bm{q})$ can be written in the form.
$\begin{split}\bm{J}_{11}(\bm{q})&=\begin{bmatrix}((\bm{p}^{Q}-\bm{p}_{1}^{J})\times\bm{\xi}_{1}(\bm{q}))^{T}\\ ((\bm{p}^{Q}-\bm{p}_{2}^{J})\times\bm{\xi}_{2}(\bm{q}))^{T}\\ ((\bm{p}^{Q}-\bm{p}_{3}^{J})\times\bm{\xi}_{3}(\bm{q}))^{T}\end{bmatrix}^{T}\\ \bm{J}_{22}(\bm{q})&=\begin{bmatrix}\bm{\xi}_{4}(\bm{q})&\bm{\xi}_{5}(\bm{q})&\bm{\xi}_{6}(\bm{q})\end{bmatrix}\end{split}$ (9)
where $\bm{\xi}_{i}(\bm{q}),i=1\dots,6$ is the unit vector pointing to the rotational direction of axis $J_{i}$ and $\bm{p}_{i}^{J},i=1\dots 6$ are any random points on axis $J_{i}$. Since J4 is perpendicular to J5 and J5 is perpendicular to J6, $\bm{J}_{22}(\bm{q})=0$ only when J4 is parallel to J6, which happens whenever $q_{5}=0$, as shown in Fig. 3B. One important arm singularity is when $\bm{p}^{Q}-\bm{p}_{2}^{J}$ and $\bm{p}^{Q}-\bm{p}_{2}^{J}$ are parallel, happening when $q_{3}=0,\pm 180^{\circ}$. These are usually called elbow singularities [36].
Arm singularities forms a well-identified set of points in the reachable space $\mathcal{RS}^{Q}$ of point $Q$, so they can be suitably and intuitively avoided by the human operator. Wrist singularities, on the other hand, can be encountered everywhere in the reachable space $\mathcal{RS}^{Q}$, making it overwhelming for the human operator to intentionally avoid them [36]. Thus, in this work, we focus on finding a motion mapping to avoid wrist singularities. In section IV-C, we will discuss this singularity-avoiding motion mapping in detail.
IV-C Singularity-Avoiding Motion Mapping
We first lock J4 at $q_{4}=0$ and generate the Z-Y-X Euler angle leading to $\bm{dR}_{t}^{m}$. Let $\bm{H}$ be the mapping from the a given rotation matrix to the Z-Y-X Euler angles.
$\begin{split}\begin{bmatrix}\theta_{t}^{m}&\phi_{t}^{m}&\psi_{t}^{m}\end{bmatrix}&=\bm{H}(\bm{R}_{t}^{m})\\ \begin{bmatrix}\theta_{0}^{m}&\phi_{0}^{m}&\psi_{0}^{m}\end{bmatrix}&=\bm{H}(\bm{R}_{0}^{m})\end{split}$ (10)
Then, instead of asking the end-effector to follow the rotational translation $\bm{dR}_{t}^{m}$, we directly add the scaled Euler angles to the target J1, J5, and J6 poses.
$\begin{split}q_{1,t}&=q_{1,t-1}+\alpha_{r}(\theta_{t}^{m}-\theta_{0}^{m})\\ q^{\prime}_{5,t}&=q^{\prime}_{5,t-1}+\alpha_{r}(\phi_{t}^{m}-\phi_{0}^{m})\\ q_{6,t}&=q_{6,t-1}+\alpha_{r}(\psi_{t}^{m}-\psi_{0}^{m})\end{split}$ (11)
$q_{1,t}$ and $q_{5,t}$ have been resolved in (11) and $q_{5,t}=q^{\prime}_{5,t}+q_{2,t}+q_{3,t}$ Notice that this is different from computing $\begin{bmatrix}d\theta_{t}^{m}&d\phi_{t}^{m}&d\psi_{t}^{m}\end{bmatrix}=\bm{H}(\bm{dR}_{t}^{m})$ and adding them to $q_{1,t}$, $q^{\prime}_{5,t}$ and $q_{6,t}$.
Without loss of generality, let the local x-direction and y-direction respectively be the normalized master forward vector and master upward vector.
$\begin{split}\bm{fw}_{t}^{m}&=\cos\phi_{t}^{m}\bm{e}_{x}+\sin\phi_{t}^{m}\bm{e}_{y}\\ \bm{up}_{t}^{m}&=\bm{e}_{z}\end{split}$ (12)
where $\bm{e}_{x}$, $\bm{e}_{y}$ and $\bm{e}_{z}$ are respectively X, Y, and Z direction unit vectors in world frame. Similarly, we also define the normalized slave forward vector and slave upward vector.
$\begin{split}\bm{FW}_{t}^{s}&=\cos\phi_{t}^{s}\bm{e}_{x}+\sin\phi_{t}^{s}\bm{e}_{y}\\ \bm{UP}_{t}^{s}&=\bm{e}_{z}\end{split}$ (13)
$q_{2,t}$ and $q_{3,t}$ should satisfy
$\begin{split}(\bm{p}_{t}^{Q}-\bm{p}_{0}^{Q})\cdot\bm{FW}_{t}^{s}&=\alpha_{l}\bm{dp}_{t}^{m}\cdot\bm{fw}_{t}^{m}\\ (\bm{p}_{t}^{Q}-\bm{p}_{0}^{Q})\cdot\bm{UP}_{t}^{s}&=\alpha_{l}\bm{dp}_{t}^{m}\cdot\bm{up}_{t}^{m}\end{split}$ (14)
where $\bm{p}_{t}^{Q}$ depends on $q_{1,t}$, $q_{2,t}$ and $q_{3,t}$ while $q_{1,t}$ is known. (14) can be easily solved using trigonometry, and this process is singularity-free for most robot configurations. Finally, $q_{5,t}$ is obtained by
$q_{5,t}=q^{\prime}_{5,t}-q_{2}-q_{3}$ (15)
Since (11) depends on the local component’s relative motion, although mapping $\bm{H}$ has singularity when $\theta_{0}^{m}$ or $\theta_{t}^{m}=\pm 90^{\circ}$, the user can easily avoid this point via resetting the starting master pose $\bm{R}_{0}^{s}$. The simple additions in (11) and (15) indicates that as long as $\theta_{t}^{m}$ or $\theta_{0}^{m}$ in the local component are kept away from ±$\pm 90^{\circ}$, the slave system motion will always be smooth even in wrist singularities.
An important property of the motion mapping in (10) and (11) is that if the starting pose of the local component $\bm{R}_{0}^{m}$ aligns with the slave end-effector $\bm{R}_{0}^{s}$, then they will be pointing at the same direction at all time, $\bm{fw}_{t}^{m}=\bm{FW}_{t}^{s}$, if $\alpha_{r}=1$.
However, this singularity avoidance mapping comes with a price, that being the lose of the J4 DoF. $q_{4}$ will be maintained at zero throughout local control. Unfortunately, as argued in section IV-B, if J4 is released and $q_{5}$ is very close to zero, a random orientation-preserved linear motion with small amplitude of the local component will cause a large movement of the slave system. Therefore, the authors believe that developing a local motion mapping that preserves all 6 DoFs of the slave end-effector while maintaining smooth transitions at singularities requires new designs on the slave side. This will be discussed in detail in section VII.
V Operational Jacobian in Teleoperation
In this part, we give a more general and mathematically rigorous analysis of why the singularity-avoiding mapping can avoid wrist singularities.
Denote the generalized position of the master device with vectors $\bm{P}$, which can either be joint poses in the global component or the spatial position defined in the $SE(3)$ set. $\bm{q}$ is the generalized position of the slave system, which usually consists of the slave joint poses. We spot that relation between the master device motion $\dot{\bm{P}}\in\mathbb{R}^{n}$ and slave side motion $\dot{\bm{q}}\in\mathbb{R}^{m}$ often has the following form
$\mathcal{F}(\bm{P})\dot{\bm{P}}=\bm{J}_{o}(\bm{P},\bm{q})\dot{\bm{q}}$ (16)
where $\mathcal{F}(\bm{P})\in\mathbb{R}^{\min(m,n)\times n}$ decreases dimension to handle the case when $n>m$. We name $\bm{J}_{o}(\bm{P},\bm{q}):\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{m\times m}$ the operational Jacobian, which can be used to indicate the smoothness of the motion mapping. The configuration that causes $\bm{J}_{o}$ to have a zero singular value is called operational singularity, meaning that certain local component motion will cause indefinite joint motion in the slave system. For our local component, which is a haptic device, there is $\dot{\bm{P}}=\begin{bmatrix}\bm{v}^{o}&\bm{\omega}^{o}\end{bmatrix}$, where $\bm{v}^{o}$ and $\bm{\omega}^{o}$ are the stylus linear and angular velocities.
In direct motion mapping as shown in Section III, there are relations $v_{i}^{o}=\alpha_{l}v_{i}$ and $\omega_{i}^{o}=\alpha_{r}\omega_{i}$ from (2), indicating $\mathcal{F}^{DM}(\bm{P})=\bm{I}_{6}$ and the following relation between the operational Jacobian and manipulator Jacobian.
$\bm{J}^{DM}_{o}(\bm{P},\bm{q}_{t})=\begin{bmatrix}\alpha_{l}\bm{I}_{3}&\bm{O}\\ \bm{O}&\alpha_{r}\bm{I}_{3}\end{bmatrix}\bm{J}_{e}(\bm{q}_{t})$ (17)
Therefore, $\bm{J}^{DM}_{o}$ inherits all singularities of the manipulator Jacobian. Since $\bm{J}^{DM}_{o}$ depends only on $\bm{q}_{t}$, it is infeasible to achieve smooth transition around singularities via adjusting the starting pose $\bm{p}_{0}^{m}$.
On the other hand, for the singularity avoidance mapping in section IV-C, since $q_{4}$ is fixed at 00, we have $m=5$ and $n=6$. Its operational Jacobian and projector can be written as the following
$\begin{split}\bm{J}^{SA}_{o}(\bm{P},\bm{q})&=\begin{bmatrix}0&A_{11}&A_{12}&0&0\\ 0&A_{21}&A_{22}&0&0\\ 1&0&0&0&\cos{\theta_{t}^{m}}\\ 0&0&0&\sin{\phi_{t}^{m}}&\sin{\theta_{t}^{m}}\cos{\phi_{t}^{m}}\\ 0&0&0&-\cos{\phi_{t}^{m}}&\sin{\theta_{t}^{m}}\sin{\phi_{t}^{m}}\end{bmatrix}W\\ \mathcal{F}^{SA}(\bm{P})&=\begin{bmatrix}\cos{\phi_{t}^{m}}&\sin{\phi_{t}^{m}}&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{bmatrix}\end{split}$
$W$ is a constant matrix generated from $q_{5}^{\prime}=q_{5}+q_{2}+q_{3}$.
$W=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&1&1&1&0\\ 0&0&0&0&1\end{bmatrix}$ (18)
Letting the distance from J2 to J3 and the distance from J3 to $Q$ to be $l_{2}$ and $l_{3}$ respectively, we have
$\begin{split}A_{11}&=l_{2}\sin{q_{2}}-l_{3}\sin(q_{2}+q_{3})\\ A_{12}&=-l_{3}\sin(q_{2}+q_{3})\\ A_{21}&=l_{2}\cos{q_{2}}-l_{3}\cos(q_{2}+q_{3})\\ A_{22}&=-l_{3}\cos(q_{2}+q_{3})\end{split}$ (19)
Since $W$ and $\mathcal{F}^{SA}(\bm{P})$ always have nonzero singular values, identifying the singularities in (V) is equivalent to identifying singularties of the following two submatrices.
$\bm{A}=\begin{bmatrix}A_{11}&A_{12}\\ A_{21}&A_{22}\\ \end{bmatrix}\quad\bm{B}=\begin{bmatrix}1&0&\cos{\theta_{t}^{m}}\\ 0&\sin{\phi_{t}^{m}}&\cos{\theta_{t}^{m}}\cos{\phi_{t}^{m}}\\ 0&-\cos{\phi_{t}^{m}}&\cos{\theta_{t}^{m}}\sin{\phi_{t}^{m}}\end{bmatrix}$
$\bm{A}$ has zero singular value at the elbow singularity, which can be easily and intuitively avoided by the human operators. $\bm{B}$ is singular when $\theta_{t}^{m}=\pm 90^{\circ}$. Interestingly, $\theta_{t}^{m}$ here is the pitch angle of the local component instead of the slave end-effector. This means one can always reset the local component starting orientation $\bm{R}_{0}^{m}$ to avoid $\theta_{t}^{m}$ from reaching ±$\pm 90^{\circ}$. Therefore, singularities in $\bm{B}$ is independent of slave joint poses $\bm{q}$, showing that wrist singularities is no longer impeding the teleoperation performance.
Motion mapping (16) covers a much wider range of teleoperation systems and does not require the master and slave systems to have the same DoF. Therefore, the idea of using operational Jacobians in analyzing the teleoperation system’s robustness against manipulator singularities can also be generalized to teleoperation systems with different kinematic structures.
Figure 5: Hardware setup of the Global–Local (G–L) teleoperation interface. (A), (B): Global teleoperation subsystem using a 3-DOF joint-link chain as the master device. (C), (D): Local teleoperation subsystem using a commercial haptic device as the master component.
Figure 6: Snapshots of tasks requiring precise manipulation. (A) Taking out Nested Buckets. (B) Key Insertion. (C) SIM Card Needle Insertion. (D) Needle Threading. (E) Wire Testing with Multimeter. The supplementary video of this paper shows the whole process of these demonstrations. Examples (B), (C), (D), and (E) feature bimanual fine manipulation, where both slave arms need to maintain precise position or orientation for completion.
VI Experimental Validations
This section presents experimental results validating the performance of G-L interface leveraging the direct motion mapping (Section III) and singularity-avoiding motion mapping (Section IV). The evaluation focuses on three key aspects:
  1. 1.
    Precision Control: Using direct motion mapping, we assessed the system’s capability for high-accuracy telemanipulation through tasks with stringent precision requirements.
  2. 2.
    Workspace Coverage: Also employing direct motion mapping, we evaluated the ability to command extensive movements across the slave robot’s workspace, demonstrating effective utilization of its kinematic limits.
  3. 3.
    Orientation Dexterity and Singularity Handling: Utilizing the singularity-avoiding mapping, we demonstrated intuitive control during tasks involving significant orientation changes, specifically highlighting smooth trajectory execution during the traversal of wrist singularities.
In addition to these demonstrations, a quantitative comparison was conducted to assess the benefits of the proposed integrated Global-Local (G-L) interface. Task completion times for both fine-grained manipulation and large-range motion scenarios were measured using the full G-L interface and compared against performance when using only the global or local component in isolation. This analysis serves to quantify the advantages derived from combining these complementary components.
VI-A Implementation Details
The global components are actuated with Dynamixel XL330-M288 servomotors arranged to a desired joint-link chain mimicking the Pieper robot arms. For direct mapping, the global component has full 6 DoFs, and for local mapping, it has only 3 DoFs. Fig. 5 shows snapshots of G-L interface realization. The motion correspondence of the direct mapping and sigularity-avoiding realizations can be found in the supplementary video. The local component comprises two 3D System 6-DoF Touch devices for precise end-effector control in Cartesian space. The isolated global components in the comparison experiments has 6 DoFs.
VI-B Precise Manipulation Demo
Figure 7: Snapshots of tasks requiring large workspace. (A) Move an object from the desk to the cabinet. (B) Take a bag from the cabinet and place it on the desk. (C) Bimanual handling of oranges from the front-left of the workspace to the back-right. (D) Bimanual transferring a test tube from a wash basin to a rack. (E) Pouring the liquid from a test tube into a beaker, and then place the test tube back on the rack.
To assess the precision and dexterity afforded by the direct mapping teleoperation realization, we conducted a series of experiments involving intricate manipulation tasks. These tasks, detailed below, were selected to represent challenges commonly encountered in real-world scenarios and to highlight the system’s capabilities in addressing them. Snapshots of these tasks are shown in Fig. 6.
Figure 8: Snapshots of daily object bimanual manipulation. (A) Opening a snack box. (B) Opening bottle cap. (C) Opening chip bag.
Figure 9: Comparison of completion time between global component only, local component only, and our proposed G-L interface. The G-L interface outperforms only global component in most tasks, except in the Nested Bucket Picking. As shown in Table I, despite taking less time to complete successful trials, only global component has a significantly lower success rate in Nested Bucket Picking. In large distance motion, G-L interface significantly outperforms local component only.
  • Key Insertion: This task, known for its demanding precision requirements, involves inserting a key into a standard bronze lock. The small tolerances between the key and lock necessitate precise control of both linear and rotational motion, as even minor deviations can result in failure or damage to the lock.
  • Needle Threading: This bimanual task requires one slave arm to manipulate a fishing line through a needle eye with a radius of approximately 1.5 mm. The second slave arm then grasps the line and pulls it through, demonstrating coordinated manipulation and fine teleoperation.
  • Taking out Nested Buckets: This task evaluates the system’s ability to manipulate objects with complex geometries and contact dynamics. Using a set of four nested buckets from the Yale-CMU-Berkeley Object and Model Set [38], the slave arms must sequentially extract each bucket and place it stably on a surface, avoiding collisions or accidental drops.
  • SIM Card Needle Insertion: This task replicates a common yet delicate manipulation requiring precise alignment. The slave robot must accurately insert a SIM card needle into the small ejector hole of a mobile phone, then apply the necessary force to release the SIM card tray. This highlights the system’s capacity for fine control in constrained environments.
  • Wire Testing with Multimeter: This task assesses the system’s ability to perform precise movements in a functional context. The slave robot uses a multimeter to test the conductivity of a Dupont wire by contacting the probes to the small metallic openings on the wire’s connectors. This task underscores the system’s suitability for tasks requiring both precision and tool manipulation.
VI-C Large Linear Range Manipulation Demo
To assess the ability of the G-L interface to effectively utilize the entire workspace of the slave robot, we conducted two demonstrations involving large-scale movements spanning the workspace’s full range along the left-right, front-back, and vertical axes. These demonstrations aimed to confirm that the integration of precise manipulation capabilities within the G-L interface did not compromise its capacity for executing large-range manipulation tasks.
  • Object Handling: This demonstration involved a coordinated bimanual task in which one slave robot retrieved an orange from the far left-front corner of its workspace and transferred it to the second slave robot. The second robot then handed off the orange to a human operator located at the far right-back corner of the workspace. This task exemplifies the system’s ability to facilitate long-range object transfer and cooperative manipulation, similar to demonstrations in prior teleoperation systems [29].
  • Pick and Place from a Cabinet: This demonstration involved interacting with a tall cabinet positioned behind the slave robot. The task required retrieving specific objects from a cluttered arrangement within the cabinet and placing them in front of the robot, or conversely, placing objects from the workspace into the cabinet. This task highlights the system’s capacity for navigating and manipulating objects within complex, spatially extended environments.
TABLE I: Success rates of isolated global component and G-L interface across different tasks. Tasks are selected based on predefined failure criteria described in Section VI-B.
Task Global Component G-L Interface
Nested Bucket Picking $3/7$ $6/7$
Pencil Picking $/7$ 7 $/7$ 7
Membrane Touching $5/7$ $/7$ 7
Buckle Opening $6/7$ $/7$ 7
Handle Lifting $/7$ 7 $/7$ 7
Both tasks are completed successfully and their snapshots are shown in Fig. 7. These tasks are good examples of how global and local components can cooperate to improve user experiences. The global component ensures a large range of motion while the local component enables the operator to make fine adjustments in clustered environments.
VI-D Large Rotational Range Manipulation Demo
To assess the ability of the singularity avoidance mapping on intuitive orientation control and robustness against wrist singularities, we conducted demonstrations involving considerable end-effector orientation change and needs to cross the wrist singularities.
  • Test Tube Pick and Place: This demonstration involves picking up a flat-laying test tube from a basin, handle it over to another robot arm and place it on a rack. This task requires significant pitch angle change and passes through 156 and 456 singularities.
  • Liquid Pouring: This demonstration involves picking up a test tube from the rack, pour the liquid into a beaker and place the test tube rack. This task requires significant yaw angle change.
Both tasks are completed efficiently and their snapshots are shown in Fig. 7D and 7E. These tasks are good examples of how singularity avoidance motion mapping yields intuitive orientation control and robustness against wrist singularities.
VI-E Daily Object Manipulation
We additionally show the G-L interface can also complement bimanual teleoperation tasks involving daily objects as in Fig. 8. These tasks include picking up a plate, where one slave robot must push down one edge to lift up the other, thereby allow another slave robot to pick up the entire plate. Tasks like opening a chip bag, snack box, and water bottle are also included.
VI-F Fine-Grained Task Completion Time Comparisons
To quantitatively assess the performance improvements enabled by the G–L hierarchical teleoperation system, we conducted a comparative study involving three control conditions: (1) using the global component only, (2) using the local component only, and (3) using both the global and local components. Seven participants were recruited to perform a series of representative tasks, each completed twice under every condition. Regrasping and retrying were allowed unless the task was classified as a failure based on predefined criteria (detailed below). Failed trials were excluded from the computation of average completion time.
Prior to the experiment, all participants received a brief tutorial on system operation and were allotted one hour of practice to familiarize themselves with the interface and task procedures. For each trial, the master and slave systems were reset to identical initial configurations, and the task completion time was recorded.
Snapshots of the following tasks designed for assessment can be found in the supplementary materials. In addition, the Needle Threading and Wire Testing with Multimeter are also included. As both tasks have no failure criteria, participants made continuous attempts until successful.
  • Nested Bucket Picking: Two nested bins were placed on a desk. Participants were tasked with lifting the inner bin without disturbing the outer bin. If either bin fell off the desk during the manipulation, the task was considered to be failed.
  • Pencil Picking: Participants were tasked with picking up a pencil from a random location on a desk using the slave robot. Contacting the desk with excessive force using the robot gripper resulted in task failure. The pencil’s position was randomized for each trial, but participants were allowed to practice with a pencil beforehand.
  • Membrane Touching: A thin piece of scotch tape, simulating a delicate membrane, was placed over the opening of a small cup. Participants were instructed to bring the needle held by the slave robot into contact with the membrane without puncturing it. Penetration constituted task failure.
  • Buckle Opening: Participants were asked to open a buckle on a box to open it using the slave robot. Since this task involves close and tight contact with an object, emergency stop of the Piper robot can be triggered if the joint torque exceeds its safety limit. The task was considered failed if the emergency stop is triggered.
  • Handle Lifting: Participants rotate a laid-down box handle to an upright position and lift the box. Failure conditions are identical to the Buckle Opening task.
  • Dots Aiming: Abstracting common insertion tasks (e.g., SIM ejector, shafts), participants use the slave gripper to prick three target dots (1 mm and 3 mm radius) with a needle. The target sheet’s location is hidden during trials but available for practice. No failure criterion; continuous attempts are allowed until success.
  • Large Distance Motion: This task compares the G–L interface with a local-only teleoperation setup to evaluate efficiency in executing large-range movements. When repeatedly performing such motions, the G–L interface demonstrates a clear advantage in operational efficiency. Specifically, the G–L interface achieves a significantly shorter task completion time, averaging only $3.30\pm 0.53$ seconds, compared to $15.1\pm 1.3$ seconds using the local component alone.
Fig. 9 presents the results of the quantitative study, demonstrating that the G–L interface consistently outperforms the isolated global or local components across most tasks. While participants using all configurations were able to quickly move the slave robot’s end-effector near the target area, those operating with the global-only setup frequently struggled to perform the final fine motions required for task completion. In contrast, the local component within the G–L interface enabled smoother and more efficient fine-tuning, resulting in significantly shorter overall completion times. The isolated global component system demonstrated marginally faster completion in nested bucket picking. However, it has a considerably lower success rate comparing to the G-L interface, as elaborated in Table I.
VII Conclusion, Limitations, and Future Work
This paper presents the Global–Local (G–L) teleoperation interface, which decouples large-scale positioning and fine manipulation through a hierarchical dual-mode control framework. We implemented both direct and singularity-avoiding motion mapping strategies, and introduced the concept of an operational Jacobian to analyze the smoothness and stability of joint behavior during local control. Experimental results across a range of tasks—including precision placement, large workspace traversal, and bimanual manipulation—demonstrate that the G–L interface improves efficiency, task success rate, and user experience compared to conventional global-only or local-only systems.
Nonetheless, several limitations remain. Manual mode switching introduces operational friction and may increase cognitive load. The current singularity-avoiding mapping strategy is task-driven and heuristic, lacking general optimality or theoretical uniqueness. Additionally, the system does not yet support force–torque feedback, and has not been validated under remote or delayed network conditions. Furthermore, the G–L interface’s modular architecture supports flexible combinations of global and local master devices, making it broadly adaptable to various application scenarios beyond the current implementation.
Future work will focus on integrating automatic or learning-based switching mechanisms, increasing the system’s degrees of freedom to improve singularity robustness, and incorporating visual and force–torque feedback for closed-loop, compliant control. We also plan to extend the operational Jacobian framework to accommodate non-Pieper structures and redundant manipulators, enabling broader applicability and more adaptive mapping strategies. These developments aim to further enhance the robustness, responsiveness, and scalability of the G–L teleoperation interface for real-world deployment in industrial, surgical, and exploratory environments.

### v3

@@
-Global-Local Interface with Selective Direct and Singularity-Avoiding Motion Mapping for Intuitive Teleoperation
+Global-Local Interface for On-Demand Teleoperation
 Abstract
-This paper presents the Global–Local (G–L) Teleoperation Interface, a hierarchical framework that enhances human–robot interaction by decoupling large-scale manipulator positioning from fine end-effector manipulation.
-The global component enables efficient workspace traversal, while the local component facilitates precise and dexterous maneuvering.
-To address kinematic singularities on the slave side—particularly during local fine manipulation—we propose a singularity-avoiding motion mapping strategy that improves both stability and intuitiveness.
-Additionally, we introduce the concept of an operational Jacobian to formally analyze the smoothness of slave joint motions under local control.
-The G–L interface is realized in two variants: Direct Mapping and Singularity-Avoiding Mapping.
-Both implementations are validated through hardware experiments involving tasks that demand precision, extended linear motion, and large orientation changes.
-Results demonstrate significant improvements in task success rate, execution efficiency, and user experience compared to conventional systems relying solely on global or local control.
-These findings indicate that the G–L interface offers a robust and scalable foundation for intuitive and versatile teleoperation across complex tasks and unstructured environments.
-Index Terms:
-Teleoperation, Human-Robot Interface, Kinematic Mapping, Singularity Avoidance
-Figure 1: Comparison of teleoperation interfaces across different system configurations.
-Our proposed G-L interface (both bimanual and single-arm variants) offers broader joint-space coverage and higher precision, while retaining compatibility with complex end-effectors.
+Teleoperation is a critical method for human-robot interface, holds significant potential for enabling robotic applications in industrial and unstructured environments.
+Existing teleoperation methods have distinct strengths and limitations in flexibility, range of workspace and precision.
+To fuse these advantages, we introduce the Global-Local (G-L) Teleoperation Interface.
+This interface decouples robotic teleoperation into global behavior, which ensures the robot motion range and intuitiveness, and local behavior, which enhances human operator’s dexterity and capability for performing fine tasks.
+The G-L interface enables efficient teleoperation not only for conventional tasks like pick-and-place, but also for challenging fine manipulation and large-scale movements.
+Based on the G-L interface, we constructed a single-arm and a dual-arm teleoperation system with different remote control devices, then demonstrated tasks requiring large motion range, precise manipulation or dexterous end-effector control.
+Extensive experiments validated the user-friendliness, accuracy, and generalizability of the proposed interface.
+Figure 1: Overview of the Global-Local Teleoperation Interface.
+The master system is decomposed into the global and local component.
+The global component controls rapid large-range motions and reflect the general collision configuration of the slave system.
+The local component controls the accurate or dexterous local motions.
+Their detailed realizations, and whether they are allowed to be activated simultaneously, can be adapted based on task requirements.
 I Introduction
-Teleoperation serves as a critical human–robot interface paradigm, enabling remote robotic operation in scenarios where human manual manipulation is insufficient or unsafe, such as surgery, disaster response, and exploration in hazardous or unstructured environments [1, 2, 3, 4, 5, 6, 7, 8].
-It also plays an increasingly essential role in data-driven robotic learning, where high-quality human demonstration datas are crucial for imitation learning and multimodal policy training [9, 10, 11, 12, 13, 14, 15, 16].
-Despite extensive research, existing teleoperation frameworks exhibit fundamental trade-offs among dexterity, workspace range, and intuitiveness.
-One common approach employs six-DoF tracking of the operator’s hand using floating controllers or vision-based systems [17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
-These enable freehand dexterity but often lack physical correspondence with the slave robot’s workspace, leading to non-intuitive motion mapping and vulnerability against manipulator singularities.
-Alternatively, kinematically matched master devices replicate the structure of the slave robot using a scaled joint-link chain, or in some cases, by employing an identical slave device as the master itself [27, 28, 29, 30, 31].
-This improves mapping of reachable configurations and collision awareness.
-However, such systems often constrain the operator’s movements and suffer from encoder detent torque, mechanical friction, and limited fidelity during fine manipulation.
-To overcome these limitations, we propose the Global–Local (G–L) Teleoperation Interface (Fig.
-1), a framework that hierarchically decouples human input into two synergistic control pathways: a global controller for large-scale positioning and a local controller for high-precision end-effector manipulation.
-This modular design fuses joint space awareness and dexterous control while minimizing the drawbacks of traditional systems.
-The G–L interface is a general teleoperation framework that supports flexible combinations of global and local master devices, allowing system configurations to be tailored to diverse tasks and hardware platforms.
-In addition, to facilitate more intuitive orientation control and avoid the impact of manipulator singularity while the local component is activated, we introduce a practical singularity-avoiding motion mapping strategy targeting the Pieper kinematics structure [32] between the local component motion and slave system’s joints.
-A concept termed operational Jacobian is proposed to give a rigorous mathematical argument of why this singularity-avoiding motion mapping ensures smooth slave joint motions at manipulator singularities.
-This concept can also be generalized to analyze the teleoperation control on slave systems other than Pieper structures.
-Experimental validation and demonstration of the G-L interface utilized two different implementations defined by the mapping from the local component to the slave system’s joints.
-The first, designated Direct Mapping, maps the scaled Cartesian motion of the local component directly to the slave end-effector.
-This preserves six degrees of freedom for local manipulation but is characterized by less intuitive orientation mapping and the absence of a singularity avoidance feature.
-In contrast, the second implementation, Singularity-Avoiding Mapping, employs the aforementioned manipulator singularity avoidance technique.
-Although this method reduces the available degrees of freedom by one, it provides more intuitive orientation mapping and robustly avoids singularities within its designated workspace.
-The contributions of this work are summarized as follows.
-  1. 1.
-    Proposal of the G-L Teleoperation Interface: A hierarchical architecture that decouples large-scale positioning from fine end-effector manipulation, designed to enhance operational flexibility and task-specific performance.
-  2. 2.
-    Implementation of Two Interface Realizations: Development and demonstration of the G-L concept through two distinct hardware implementations: the Direct Mapping Realization and the Singularity-Avoiding Realization.
-The later introduces a specialized motion mapping strategy for the local control component, aimed at improving orientation control intuitiveness and ensuring stable operation near slave manipulator singularities.
-  3. 3.
-    Introduction of the Operational Jacobian: Proposal of the operational Jacobian as a mathematical tool to rigorously analyze the smoothness of the slave system’s joint motion during teleoperation.
-  4. 4.
-    Experimental Validation and Comparative Analysis: Comprehensive validation of the G-L interface through experiments involving precision manipulation, large linear movements, and significant angular adjustments.
-Results demonstrate quantifiable improvements in success rates, operational efficiency, and user experience compared to conventional single-mode (global-only or local-only) systems.
-The remainder of this paper is organized as follows.
-Section II introduces the proposed Global-Local (G-L) teleoperation interface and its key design criteria.
-Section III details the direct mapping realization, while Section IV presents the singularity-avoiding mapping strategy.
-Section V describes the operational Jacobian tool.
-Experimental validations across various manipulation tasks are discussed in Section VI.
-Finally, Section VII concludes the paper with a summary of contributions, limitations, and directions for future work.
+Feature G-L Interface (Bi) G-L Interface (Si) Gello [1] ACE[2] OpenTelevision [3]
+End-effector Type Two-fingered Gripper Dexterous Hand Two-Fingered Gripper Gripper and dexterous hand Dexterous hand
+Arm Mode Bimanual Single Bimanual Bimanual Bimanual
+Master Device Arm replica + Touch X Arm replica + IMU + exoskeleton Arm Replica Arm replica + RGBD camera VR (VisionPro)
+Teleoperation of Arm Joint mapping + Inverse Kinematics Joint mapping + Euler angle Joint mapping Joint mapping Inverse kinematics
+Teleoperation of End-effector Boolean button Joint mapping Joint mapping Hand tracking Hand tracking
+Motion Range Full joint space Full joint space Partial joint space Front cartesian space Front cartesian space
+In-device collision None None Restricted None None
+Precision High Mid Mid Mid Mid
+TABLE I: Comparison of Teleoperation Interfaces
+With the advancement of embodied intelligence, there is an increasing expectation for smarter robots capable of operating across a broader range of scenarios, including industrial settings and unstructured environments [4, 5].
+Teleoperation has gained growing recognition as a critical human-robot interface [6, 7, 8].
+Teleoperation has demonstrated exceptional value across various domains and scenarios, such as robotic surgery, where precise movements and real-time decision-making are indispensable [9, 10, 11]; underwater exploration, which grants access to hazardous and challenging environments [12]; space exploration, extending human reach beyond Earth’s boundaries for scientific discoveries and exploration [13, 14]; and disaster response, providing essential assistance in dangerous situations, such as search and rescue operations [15, 16, 17].
+In addition, teleoperation offers efficient and cost-effective means for collecting high-quality demonstration data [1, 7].
+This is crucial for advancing robotic learning frameworks, such as imitation learning [18, 19, 20] and vision-language-action (VLA) models [21, 22], where data availability has become a significant bottleneck [23, 24, 25].
+Although various teleoperation approaches have been proposed, each exhibits distinct strengths and limitations.
+Some typical methods utilize a remote floating device, representing the master’s end-effector, and an external tracking system, often vision-based, to determine its pose [26, 27, 28, 3, 29, 30, 31, 32, 2, 24].
+Under this framework, the master system almost always has six degrees of freedom (DoF), representing the operator’s hand position and orientation.
+While this approach facilitates greater operator dexterity due to the absence of mechanical constraints, challenges arise in both aligning the workspaces of the master and slave systems and mitigating the loss in the super-human’s working space intrinsic to the robot arm.
+Even when the master and slave possess the same number of DoF, discrepancies in their workspace geometries can lead to situations where the operator specifies unattainable poses for the slave, or difficulty for the operator to drive the slave system to a reachable point.
+This issue is further compounded by differences in the number of DoF.
+When the slave possesses fewer than six DoF, the likelihood of specifying unattainable poses increases.
+Conversely, for slaves with more than six DoF, the framework cannot adequately control the slave’s nullspace, hindering the operator’s capability to leverage redundant DoF for more complicated tasks such as collision avoidance.
+On the other hand, employing a scaled replica of the slave system as the master device is also a popular design.
+This framework provides the operator with an intuitive understanding of the slave robot’s workspace, including reachability and potential collision configurations, due to the preserved kinematic structure [33, 34, 1, 18, 35, 36].
+However, the inherent mechanical constraints, teleoperation redundancy, and detent torque in encoders of the master system can impose unwanted movements or apply varying external forces which can significantly hinder the operator’s dexterity and capability to perform fine-grained tasks.
+Recognizing the limitations of existing teleoperation frameworks, we propose the novel Global-Local (G-L) Teleoperation Interface.
+The G-L interface addresses teleoperation challenges by decoupling the master system into global and local components.
+The global component facilitates efficient navigation and reflects the collision configurations of the slave system within its workspace.
+This component enables rapid, large-scale adjustments of the slave’s pose, ensuring comprehensive workspace coverage.
+The local component focuses on fine-grained control of the slave’s end-effector, delivering the dexterity and precision required for intricate tasks.
+By integrating these complementary components, the G-L Teleoperation Interface fully leverages the slave system’s capabilities, overcoming human limitations while preserving the operator’s precision and dexterity, and ultimately enables the execution of complex tasks requiring both broad reach and fine manipulation.
+Fig 1 shows a schematic diagram of how G-L interface operates.
+The experiments demonstrate the enhanced dexterity, precision, and motion range of two specialized task-specific teleoperation systems under the G-L interface.
 II Global-Local Teleoperation Interface
+This section details the decoupling of master systems into global and local components within the G-L Teleoperation Interface.
 To be considered a valid implementation of the G-L Interface, a decoupled master system must adhere to the following criteria.
@@
 Second, the local component must facilitate precise, fine-grained control of the slave’s end-effector, enabling the operator to execute tasks requiring dexterity and accuracy.
-It should prioritize the seamless transfer of the operator’s manipulation skills to the slave system, with less emphasis on reflecting the slave’s overall configuration or controlling all DoFs.
-In the remaining of this paper, without further clarification, our discussions are limited to slave systems with serial connected structure.
-Employing a scaled replica of the slave system, a common strategy in teleoperation [9, 30], effectively meets the requirements for the global component of our system.
-However, this implementation differs from existing scaled replica teleoperation systems in two ways.
-Firstly, the global component does not require exact kinematic correspondence with the slave system’s revolute axes, as its primary function is to enable rapid and coarse motions.
-It only needs to provide an approximation of the current slave system’s collision configuration, allowing the human operator to avoid potential collisions during coarse adjustments of the slave’s pose.
-Secondly, the global component need not encompass all Degrees of Freedom (DoFs) of the slave system.
-It is generally observed that DoFs closer to the base exert a more substantial influence on both the robot’s end-effector position and its overall collision state compared to DoFs nearer the end-effector.
-Consequently, it may be sufficient for the global component to replicate only these more influential, base-proximal DoFs.
-Tracking the operator’s hand motion represents a suitable approach for implementing the local component of the interface.
-A key challenge in Global-Local (G-L) interface design is establishing an intuitive motion mapping between this local component and the slave system.
-Focusing on the widely applied Pieper kinematics manipulator structure, we propose two distinct motion mapping methods.
-The first, more direct method involves mapping the local component’s motion to the slave system’s end-effector motion and solving for the required joint poses using inverse kinematics.
-Owing to its straightforward implementation and preservation of all 6 DoFs for the slave end-effector, this technique is commonly adopted in robot data collection applications [25].
-However, if the initial orientation of the local component does not align well with the slave end-effector’s starting orientation, this method can prove counter-intuitive for orientation control.
-Furthermore, manipulator singularities can adversely affect the usability of the local component, an issue detailed further in Section IV.
-The second method maps the three rotational DoFs of the local component to three mutually perpendicular slave joints and locks one slave DoF to mitigate certain rotational singularities.
-While this approach reduces the effective mapped DoFs by one, it provides considerably more intuitive orientation control.
-Using the two local motion mapping approaches and with different number of DoFs the global components cover, we present the Direct Mapping and Singularity-Avoiding Mapping realizations of the G-L interface, which will be introduced in detail in sections III and IV.
-Both realizations share a challenge during the transition from the local to global component.
+It should prioritize the seamless transfer of the operator’s manipulation skills to the slave system, with less emphasis on reflecting the slave’s overall configuration or controlling all degrees of freedom (DOFs).
+Finally, the combined functionality of the global and local components must ensure complete control over all DOFs of the slave system.
+This work mainly explores two decoupling methods for realizing the G-L interface which we have termed as temporal and spatial decoupling.
+In temporal decoupling, both global and local components possess full control over all slave DOFs.
+However, they are activated sequentially.
+The operator first utilizes the global component for gross positioning of the slave system, followed by the local component for fine adjustments to achieve the desired task.
+Spatial decoupling divides the slave DOFs into two groups.
+The global component controls the first group, primarily influencing end-effector position and overall slave collision configuration.
+The local component governs the second group, primarily determining end-effector orientation with minor impact on the overall configuration.
+The global and local components can be used simultaneously since they control different DOFs.
+It is important to note that these methods are not exhaustive.
+The G-L interface accommodates various decoupling strategies tailored to specific slave systems and task demands.
+The subsequent subsections elaborate on the design and implementation of temporal and spatial decoupling. Section III will showcase their physical realizations.
+II-A Temporal Decoupling
+In the temporal decoupling scheme, the global component is realized as a scaled replica of the slave system, a common approach in teleoperation [18, 35].
+The local component utilizes a haptic device for precise tracking of the operator’s fingertip position and orientation.
+Activating the global component places the slave system in joint control mode, replicating the scaled replica’s joint positions within the limits of the motor capacities.
+Conversely, activating the local component switches the slave system to end-effector Cartesian control, where it tracks the motion of the haptic device.
+Fig.
+2 shows a schematic diagram of temporal decoupling.
+To facilitate precise control of the end-effector, the slave system is designed to track only a portion of the local component motion.
+Every time the local component is activated, its current position $p_{0}$ and orientation $q_{0}$ will be re-initialized as the origin.
+Given a linear scaling factor $\alpha_{l}\in(0,1]$, the slave system will move its end-effector position towards the scaled linear displacement of the haptic device, $\alpha_{l}(p-p_{0})$.
+For rotational motions, assume the orientation of the local component has moved from $q_{0}$ to $q$, we convert the rotational displacement $q_{0}^{-1}q$ into axis-angle representation $(\vec{v},\theta)$, where $\vec{v}$ is the rotational axis and $\theta$ is the rotational angle.
+Assume the slave system’s end-effector currently has orientation $q_{e}$, given a rotational scaling factor $\alpha_{r}\in(0,1]$, the slave system will target to change its end-effector orientation via rotating around axis $\vec{v}$ by angle $\alpha_{r}\theta$, starting from $q_{e}$.
+Figure 2: A schematic diagram of temporal decoupling.
+The global component is a scaled replica of the slave robot and the local component is a haptic device.
+The two step pedals allows the operator to rapidly switch between the local and global component.
+If the local component is activated, the global component will temporary turn into a slave system that follows the actual slave system to avoid joint position mismatch when switching back to the global component.
+This design presents a challenge during the transition from the local to global component.
 A mismatch can arise between the pose of the scaled replica (global component) and the actual slave system configuration, potentially causing an abrupt motion in the slave upon returning to global control.
@@
 This ensures pose synchronization between the global component and the slave, facilitating a smooth transition when switching back to global controls.
+It is worth noting that there is an overlap between the DOFs that the global and local components are controlling, so they cannot be activated at the same time.
+Using devices like foot pedals allows the operator to conveniently switch between the global and local components.
+II-B Spatial Decoupling
+Figure 3: A schematic diagram of spatial decoupling.
+The global component is a scaled replcia of the first N3N-3 DOFs, while the local component consists of two IMU sensors to measure human operator’s wrist rotation and map it back to the last three joints’s target positions.
+Since it eliminates all mechanical constraints around the operator’s hand, it allows mounting complicated end-effector controllers, such as an exoskeleton structure for controlling a dexterous hand.
+Spatial decoupling leverages a common characteristic of serially-linked industrial robots in that the final three joints primarily influence end-effector orientation with less impact on the overall robot configuration, while the remaining joints predominantly determine end-effector position and potential collisions.
+This principle is applied to an $N$-DOF industrial robot (e.g., KUKA or FANUC) by dividing control into two components.
+The global component replicates the first N3N-3 joints, providing gross positioning, while a local component, consisting of a dual-IMU system, mounted on the forearm and the hand of the human operator.
+measures the relative rotation $R$ between the human wrist and hand to control the final three joints.
+This approach often simplifies inverse kinematics to finding X-Y-Z or X-Y-X Euler angles based on $R$.
 Fig.
-2 illustrates a representative configuration of the proposed Global–Local (G–L) interface.
-Haptic devices are used to capture the operator’s fine hand motions, while a scaled-down 3-DOF joint-link chain captures the large range arm movement.
-Two pairs of foot pedals allow the operator to switch between global and local control modes.
-Figure 2: A schematic diagram of G-L interface.
-The global component is a scaled replica of the slave robot and the local component is a haptic device.
-The two step pedals allows the operator to rapidly switch between the local and global component.
-If the local component is activated, the global component will temporary turn into a slave system that follows the actual slave system to avoid joint position mismatch when switching back to the global component.
-III Direct Mapping of Local Component Motion
-It is important to emphasize the key distinction between the Direct Mapping and Singularity-Avoiding Mapping strategies: they primarily differ in the local motion mapping method and the number of degrees of freedom allocated to the global component.
-The global component of the Direct Mapping implementation covers the full six DoFs in the Pieper structure slave robot [33, 34, 35].
-Activating the global component places the slave system in joint control mode, replicating the scaled replica’s joint positions within the limits of the motor capacities.
-Conversely, activating the local component switches the slave system to end-effector Cartesian control, where it tracks the motion of the haptic device.
-To facilitate precise control of the end-effector, the slave system is designed to track only a portion of the local component motion.
-Denote the slave end-effector pose with $(\bm{p},\bm{R})$ where $\bm{p}\in\mathbb{R}^{3}$ is the position and $\bm{R}\in\mathbb{R}^{3\times 3}$ is the orientation represented by a rotational matrix.
-Let $\bm{q}$ be the joint poses of the slave robot, then we have $\bm{q}=\bm{F}(\bm{p},\bm{R})$ where $\bm{F}$ is the inverse kinematics of Pieper kinematics structure, which can be efficiently computed in closed form [32].
-Every time the local component is activated, its starting pose $(\bm{p}^{m}_{0},\bm{R}^{m}_{0})$ as well as the slave end-effector’s starting pose $(\bm{p}^{s}_{0},\bm{R}^{s}_{0})$ will be recorded.
-Suppose at a certain moment $t$, the local component moves to pose $(\bm{p}^{m}_{t},\bm{R}^{m}_{t})$, we first compute the linear and rotational displacements since activation.
-$\begin{split}\bm{dp}_{t}^{m}&=\bm{p}_{t}^{m}-\bm{p}_{0}^{m}\\ (\bm{l}^{m}_{t},\theta^{m}_{t})&=\bm{G}(\bm{dR}^{m}_{t})=\bm{G}(\bm{R}^{m,T}_{0}R^{m}_{t})\end{split}$ (1)
-where the superscript $T$ stands for transpose and $\bm{G}$ maps the rotational matrix to axis-angle format. $\bm{l}_{t}$ is the normalized rotational axis direction and $\theta_{t}$ is the angle.
-Given prespecified linear and rotational scaling factors $\alpha_{l},\alpha_{r}\in(0,1]$, the slave system will move its end-effector pose towards the scaled linear and rotational displacements of the haptic device.
-Mathematically, the target slave pose at moment $t$ is
-$\begin{split}\bm{p}_{t}^{s}&=\bm{p}_{0}^{s}+\alpha_{l}\bm{dp}_{t}^{m}\\ \bm{R}_{t}^{s}&=\bm{R}_{0}^{s}\bm{G}^{-1}(\bm{l}^{m}_{t},\alpha_{r}\theta^{m}_{t})\end{split}$ (2)
-with $\bm{G}^{-1}$ being the mapping that gets the rotational matrix from the axis-angle interpretation.
-With the generated end-effector pose, the target joint poses of the slave robot at moment $t$ will be
-$\bm{q}^{s}_{t}=\bm{F}(\bm{p}^{s}_{t},\bm{R}^{s}_{t})$ (3)
-IV Singularity-Avoiding Mapping of Local Component Motion
-Despite the preservation of full degrees of freedom of the end-effector, the aforementioned direct motion mapping has two major limitations.
-  • It is particularly vulnerable against manipulator singularities, especially the wrist singularity in Pieper kinematics, which can happen everywhere inside the manipulator reachable workspace [36].
-This demands human operators to pay extra attention on the slave joint poses other than the end-effector pose.
-  • When the starting orientation of the local component and the slave end-effector have an obvious misalignment, controlling orientation using direct mapping can be overwhelming for human operators.
-This is because it is difficult for human operators to predict the slave end-effector’s orientation change given the motion of the local component.
-Additionally, they cannot intuitively discern which joints will be actuated and how much they will move.
-This section focuses on the scenario when the slave system has a Pieper kinematic structure and proposes a practical local component motion mapping method to alleviate these limitations.
-IV-A Pieper Kinematics and Translation-Rotation Decoupling
-Pieper kinematics structure, also known as the 321 kinematics structure, is widely adopted in 6-joint industrial robot arms for its closed-form inverse kinematics.
-Fig.
-3 is an illustration of the three most important characteristics of Pieper kinematics structure using AgileX Piper as an example.
-  • The base joint (J1) and the first shoulder joint (J2) intersect and are perpendicular.
-  • The two shoulder joints (J2 and J3) are parallel.
-  • The three wrist joints (J4, J5 and J6) intersects at the same point $Q$, usually arranged in an intrinsic X-Y-X Euler Angles order.
-Figure 3: (A) Pieper Kinematics Structure using Agile X Piper as an example.
-(B) $q_{5}=0^{\circ}$ wrist singularity, which is a typical manipulation pose and is also the singularity type we are focusing on avoiding.
-It can be observed that the intersection point $Q$ is relatively close to the end-effector tip and its position is only affected by the poses of the base and shoulder joints.
-In addition, since the three wrist joints formulate an Euler angle combinations, they can dexterously adjust the end-effector orientation in a nonzero-measure set without actuating J1, J2 and J3, regardless of the Cartesian position of $Q$ [37].
-With this observation, the following amendments are made to the Direct Mapping realization.
-  • Translation of the local component now controls the linear motion of the intersection point $Q$, $\bm{p}_{t}^{Q}$, not the end-effector’s linear motion $\bm{p}_{t}^{s}$.
-Therefore, only J1, J2 and J3 will be actuated.
-  • Rotation of the local component still controls the slave end-effector’s orientation $\bm{R}_{t}^{s}$, but only J4, J5 and J6 will be actuated.
-Denote the starting pose of point $Q$ as $(\bm{p}_{0}^{Q},\bm{R}_{0}^{Q})$, the target position of point $Q$ at moment $t$ will be
-$\bm{p}_{t}^{Q}=\bm{p}_{0}^{Q}+\alpha_{l}\bm{dp}_{t}^{m}$ (4)
-The joint poses are now solved as
-$\begin{split}\begin{bmatrix}q_{1,t}^{s},q_{2,t}^{s},q_{3,t}^{s}\end{bmatrix}^{T}&=\bm{F}_{123}(\bm{p}_{t}^{Q})\\ \begin{bmatrix}q_{4,t}^{s},q_{5,t}^{s},q_{6,t}^{s}\end{bmatrix}^{T}&=\bm{F}_{456}(\bm{R}_{t}^{Q,T}\bm{R}_{t}^{s})\end{split}$ (5)
-where $\bm{F}_{123}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ is the mapping from the $Q$ point position to J1, J2 and J3 joint poses.
-$\bm{F}_{456}:\mathbb{R}^{3\times 3}\rightarrow\mathbb{R}^{3}$ maps from rotation matrix to J4, J5 and J6 joint poses.
-$\bm{R}_{t}^{Q}$ is the current orientation of point $Q$.
-This decoupling process of the local component’s translation and rotation partially reduces the corresponding number of joints of each local component DoF to less than or equal to three.
-Although this increases control intuitiveness for human operators, it still suffers from manipulator singularities.
-In section IV-B, we discuss further amendments that can help avoid manipulator singularities.
-Figure 4: Visualization of Singularity Avoidance Motion Mapping.
-The digital pen in the center indicates the direction where the local component is pointing to.
-(A) Row angle $d\theta_{t}^{m}$ maps to J1 motion.
-Notice that under motion mapping (11), the digital pen does not necessaily align with the X-axis at the beginning.
-(B) Pitch angle $d\phi_{t}^{m}$ maps to J5 motion.
-(C) Yaw angle $d\psi_{t}^{m}$ maps to J6 motion.
-IV-B Singularities in Pieper Kinematics
-Following notations in [36], the motion mapping from the configuration space to the operational space has the following form.
-$\bm{v}^{e}_{t}=\bm{J}(\bm{q}_{t})\dot{\bm{q}_{t}}$ (6)
-An important difference here is that $\bm{v}^{e}_{t}$ now consists of the linear velocity of point $Q$ and the angular velocity of the end-effector $\bm{v}^{e}_{t}=\begin{bmatrix}\dot{\bm{p}}^{Q}_{t}&\bm{\omega}^{e}_{t}\end{bmatrix}$.
-Since Pieper kinematics structure has six independent joints, the Jacobian can be naturally partitioned into $(3\times 3)$ blocks.
-$\bm{J}(\bm{q}_{t})=\begin{bmatrix}\bm{J}_{11}(\bm{q}_{t})&\bm{J}_{12}(\bm{q}_{t})\\ \bm{J}_{21}(\bm{q}_{t})&\bm{J}_{22}(\bm{q}_{t})\end{bmatrix}$ (7)
-Manipulator singularities are the joint configurations $\bm{q}$ where the matrix $\bm{J}(\bm{q})$ presents a zero singular value.
-When $\bm{J}(\bm{q})$ is a square matrix, this is equivalent to having a zero determinant, $\det{\bm{J}(q)}=0$.
-Since the position $\bm{p}_{t}^{Q}$ is independent of J4-J6, we know that $\bm{J}_{12}(\bm{q}_{t})=\bm{O}_{3}$.
-As a result, manipulator singularities can be identified from the following equation.
-$\det\bm{J}_{11}(\bm{q})\det\bm{J}_{22}(\bm{q})=0$ (8)
-Scenarios when $\det\bm{J}_{11}(\bm{q})=0$ is often referred to as arm singularities since it mainly affects the linear velocity and $\det\bm{J}_{22}(\bm{q})=0$ is often referred to as wrist singularities since it mainly affect angular velocities.
-$\bm{J}_{11}(\bm{q})$ and $\bm{J}_{22}(\bm{q})$ can be written in the form.
-$\begin{split}\bm{J}_{11}(\bm{q})&=\begin{bmatrix}((\bm{p}^{Q}-\bm{p}_{1}^{J})\times\bm{\xi}_{1}(\bm{q}))^{T}\\ ((\bm{p}^{Q}-\bm{p}_{2}^{J})\times\bm{\xi}_{2}(\bm{q}))^{T}\\ ((\bm{p}^{Q}-\bm{p}_{3}^{J})\times\bm{\xi}_{3}(\bm{q}))^{T}\end{bmatrix}^{T}\\ \bm{J}_{22}(\bm{q})&=\begin{bmatrix}\bm{\xi}_{4}(\bm{q})&\bm{\xi}_{5}(\bm{q})&\bm{\xi}_{6}(\bm{q})\end{bmatrix}\end{split}$ (9)
-where $\bm{\xi}_{i}(\bm{q}),i=1\dots,6$ is the unit vector pointing to the rotational direction of axis $J_{i}$ and $\bm{p}_{i}^{J},i=1\dots 6$ are any random points on axis $J_{i}$.
-Since J4 is perpendicular to J5 and J5 is perpendicular to J6, $\bm{J}_{22}(\bm{q})=0$ only when J4 is parallel to J6, which happens whenever $q_{5}=0$, as shown in Fig.
-3B.
-One important arm singularity is when $\bm{p}^{Q}-\bm{p}_{2}^{J}$ and $\bm{p}^{Q}-\bm{p}_{2}^{J}$ are parallel, happening when $q_{3}=0,\pm 180^{\circ}$.
-These are usually called elbow singularities [36].
-Arm singularities forms a well-identified set of points in the reachable space $\mathcal{RS}^{Q}$ of point $Q$, so they can be suitably and intuitively avoided by the human operator.
-Wrist singularities, on the other hand, can be encountered everywhere in the reachable space $\mathcal{RS}^{Q}$, making it overwhelming for the human operator to intentionally avoid them [36].
-Thus, in this work, we focus on finding a motion mapping to avoid wrist singularities.
-In section IV-C, we will discuss this singularity-avoiding motion mapping in detail.
-IV-C Singularity-Avoiding Motion Mapping
-We first lock J4 at $q_{4}=0$ and generate the Z-Y-X Euler angle leading to $\bm{dR}_{t}^{m}$. Let $\bm{H}$ be the mapping from the a given rotation matrix to the Z-Y-X Euler angles.
-$\begin{split}\begin{bmatrix}\theta_{t}^{m}&\phi_{t}^{m}&\psi_{t}^{m}\end{bmatrix}&=\bm{H}(\bm{R}_{t}^{m})\\ \begin{bmatrix}\theta_{0}^{m}&\phi_{0}^{m}&\psi_{0}^{m}\end{bmatrix}&=\bm{H}(\bm{R}_{0}^{m})\end{split}$ (10)
-Then, instead of asking the end-effector to follow the rotational translation $\bm{dR}_{t}^{m}$, we directly add the scaled Euler angles to the target J1, J5, and J6 poses.
-$\begin{split}q_{1,t}&=q_{1,t-1}+\alpha_{r}(\theta_{t}^{m}-\theta_{0}^{m})\\ q^{\prime}_{5,t}&=q^{\prime}_{5,t-1}+\alpha_{r}(\phi_{t}^{m}-\phi_{0}^{m})\\ q_{6,t}&=q_{6,t-1}+\alpha_{r}(\psi_{t}^{m}-\psi_{0}^{m})\end{split}$ (11)
-$q_{1,t}$ and $q_{5,t}$ have been resolved in (11) and $q_{5,t}=q^{\prime}_{5,t}+q_{2,t}+q_{3,t}$ Notice that this is different from computing $\begin{bmatrix}d\theta_{t}^{m}&d\phi_{t}^{m}&d\psi_{t}^{m}\end{bmatrix}=\bm{H}(\bm{dR}_{t}^{m})$ and adding them to $q_{1,t}$, $q^{\prime}_{5,t}$ and $q_{6,t}$.
-Without loss of generality, let the local x-direction and y-direction respectively be the normalized master forward vector and master upward vector.
-$\begin{split}\bm{fw}_{t}^{m}&=\cos\phi_{t}^{m}\bm{e}_{x}+\sin\phi_{t}^{m}\bm{e}_{y}\\ \bm{up}_{t}^{m}&=\bm{e}_{z}\end{split}$ (12)
-where $\bm{e}_{x}$, $\bm{e}_{y}$ and $\bm{e}_{z}$ are respectively X, Y, and Z direction unit vectors in world frame.
-Similarly, we also define the normalized slave forward vector and slave upward vector.
-$\begin{split}\bm{FW}_{t}^{s}&=\cos\phi_{t}^{s}\bm{e}_{x}+\sin\phi_{t}^{s}\bm{e}_{y}\\ \bm{UP}_{t}^{s}&=\bm{e}_{z}\end{split}$ (13)
-$q_{2,t}$ and $q_{3,t}$ should satisfy
-$\begin{split}(\bm{p}_{t}^{Q}-\bm{p}_{0}^{Q})\cdot\bm{FW}_{t}^{s}&=\alpha_{l}\bm{dp}_{t}^{m}\cdot\bm{fw}_{t}^{m}\\ (\bm{p}_{t}^{Q}-\bm{p}_{0}^{Q})\cdot\bm{UP}_{t}^{s}&=\alpha_{l}\bm{dp}_{t}^{m}\cdot\bm{up}_{t}^{m}\end{split}$ (14)
-where $\bm{p}_{t}^{Q}$ depends on $q_{1,t}$, $q_{2,t}$ and $q_{3,t}$ while $q_{1,t}$ is known.
-(14) can be easily solved using trigonometry, and this process is singularity-free for most robot configurations.
-Finally, $q_{5,t}$ is obtained by
-$q_{5,t}=q^{\prime}_{5,t}-q_{2}-q_{3}$ (15)
-Since (11) depends on the local component’s relative motion, although mapping $\bm{H}$ has singularity when $\theta_{0}^{m}$ or $\theta_{t}^{m}=\pm 90^{\circ}$, the user can easily avoid this point via resetting the starting master pose $\bm{R}_{0}^{s}$.
-The simple additions in (11) and (15) indicates that as long as $\theta_{t}^{m}$ or $\theta_{0}^{m}$ in the local component are kept away from ±$\pm 90^{\circ}$, the slave system motion will always be smooth even in wrist singularities.
-An important property of the motion mapping in (10) and (11) is that if the starting pose of the local component $\bm{R}_{0}^{m}$ aligns with the slave end-effector $\bm{R}_{0}^{s}$, then they will be pointing at the same direction at all time, $\bm{fw}_{t}^{m}=\bm{FW}_{t}^{s}$, if $\alpha_{r}=1$.
-However, this singularity avoidance mapping comes with a price, that being the lose of the J4 DoF.
-$q_{4}$ will be maintained at zero throughout local control.
-Unfortunately, as argued in section IV-B, if J4 is released and $q_{5}$ is very close to zero, a random orientation-preserved linear motion with small amplitude of the local component will cause a large movement of the slave system.
-Therefore, the authors believe that developing a local motion mapping that preserves all 6 DoFs of the slave end-effector while maintaining smooth transitions at singularities requires new designs on the slave side.
-This will be discussed in detail in section VII.
-V Operational Jacobian in Teleoperation
-In this part, we give a more general and mathematically rigorous analysis of why the singularity-avoiding mapping can avoid wrist singularities.
-Denote the generalized position of the master device with vectors $\bm{P}$, which can either be joint poses in the global component or the spatial position defined in the $SE(3)$ set.
-$\bm{q}$ is the generalized position of the slave system, which usually consists of the slave joint poses.
-We spot that relation between the master device motion $\dot{\bm{P}}\in\mathbb{R}^{n}$ and slave side motion $\dot{\bm{q}}\in\mathbb{R}^{m}$ often has the following form
-$\mathcal{F}(\bm{P})\dot{\bm{P}}=\bm{J}_{o}(\bm{P},\bm{q})\dot{\bm{q}}$ (16)
-where $\mathcal{F}(\bm{P})\in\mathbb{R}^{\min(m,n)\times n}$ decreases dimension to handle the case when $n>m$.
-We name $\bm{J}_{o}(\bm{P},\bm{q}):\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{m\times m}$ the operational Jacobian, which can be used to indicate the smoothness of the motion mapping.
-The configuration that causes $\bm{J}_{o}$ to have a zero singular value is called operational singularity, meaning that certain local component motion will cause indefinite joint motion in the slave system.
-For our local component, which is a haptic device, there is $\dot{\bm{P}}=\begin{bmatrix}\bm{v}^{o}&\bm{\omega}^{o}\end{bmatrix}$, where $\bm{v}^{o}$ and $\bm{\omega}^{o}$ are the stylus linear and angular velocities.
-In direct motion mapping as shown in Section III, there are relations $v_{i}^{o}=\alpha_{l}v_{i}$ and $\omega_{i}^{o}=\alpha_{r}\omega_{i}$ from (2), indicating $\mathcal{F}^{DM}(\bm{P})=\bm{I}_{6}$ and the following relation between the operational Jacobian and manipulator Jacobian.
-$\bm{J}^{DM}_{o}(\bm{P},\bm{q}_{t})=\begin{bmatrix}\alpha_{l}\bm{I}_{3}&\bm{O}\\ \bm{O}&\alpha_{r}\bm{I}_{3}\end{bmatrix}\bm{J}_{e}(\bm{q}_{t})$ (17)
-Therefore, $\bm{J}^{DM}_{o}$ inherits all singularities of the manipulator Jacobian.
-Since $\bm{J}^{DM}_{o}$ depends only on $\bm{q}_{t}$, it is infeasible to achieve smooth transition around singularities via adjusting the starting pose $\bm{p}_{0}^{m}$.
-On the other hand, for the singularity avoidance mapping in section IV-C, since $q_{4}$ is fixed at 00, we have $m=5$ and $n=6$. Its operational Jacobian and projector can be written as the following
-$\begin{split}\bm{J}^{SA}_{o}(\bm{P},\bm{q})&=\begin{bmatrix}0&A_{11}&A_{12}&0&0\\ 0&A_{21}&A_{22}&0&0\\ 1&0&0&0&\cos{\theta_{t}^{m}}\\ 0&0&0&\sin{\phi_{t}^{m}}&\sin{\theta_{t}^{m}}\cos{\phi_{t}^{m}}\\ 0&0&0&-\cos{\phi_{t}^{m}}&\sin{\theta_{t}^{m}}\sin{\phi_{t}^{m}}\end{bmatrix}W\\ \mathcal{F}^{SA}(\bm{P})&=\begin{bmatrix}\cos{\phi_{t}^{m}}&\sin{\phi_{t}^{m}}&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{bmatrix}\end{split}$
-$W$ is a constant matrix generated from $q_{5}^{\prime}=q_{5}+q_{2}+q_{3}$.
-$W=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&1&1&1&0\\ 0&0&0&0&1\end{bmatrix}$ (18)
-Letting the distance from J2 to J3 and the distance from J3 to $Q$ to be $l_{2}$ and $l_{3}$ respectively, we have
-$\begin{split}A_{11}&=l_{2}\sin{q_{2}}-l_{3}\sin(q_{2}+q_{3})\\ A_{12}&=-l_{3}\sin(q_{2}+q_{3})\\ A_{21}&=l_{2}\cos{q_{2}}-l_{3}\cos(q_{2}+q_{3})\\ A_{22}&=-l_{3}\cos(q_{2}+q_{3})\end{split}$ (19)
-Since $W$ and $\mathcal{F}^{SA}(\bm{P})$ always have nonzero singular values, identifying the singularities in (V) is equivalent to identifying singularties of the following two submatrices.
-$\bm{A}=\begin{bmatrix}A_{11}&A_{12}\\ A_{21}&A_{22}\\ \end{bmatrix}\quad\bm{B}=\begin{bmatrix}1&0&\cos{\theta_{t}^{m}}\\ 0&\sin{\phi_{t}^{m}}&\cos{\theta_{t}^{m}}\cos{\phi_{t}^{m}}\\ 0&-\cos{\phi_{t}^{m}}&\cos{\theta_{t}^{m}}\sin{\phi_{t}^{m}}\end{bmatrix}$
-$\bm{A}$ has zero singular value at the elbow singularity, which can be easily and intuitively avoided by the human operators.
-$\bm{B}$ is singular when $\theta_{t}^{m}=\pm 90^{\circ}$.
-Interestingly, $\theta_{t}^{m}$ here is the pitch angle of the local component instead of the slave end-effector.
-This means one can always reset the local component starting orientation $\bm{R}_{0}^{m}$ to avoid $\theta_{t}^{m}$ from reaching ±$\pm 90^{\circ}$.
-Therefore, singularities in $\bm{B}$ is independent of slave joint poses $\bm{q}$, showing that wrist singularities is no longer impeding the teleoperation performance.
-Motion mapping (16) covers a much wider range of teleoperation systems and does not require the master and slave systems to have the same DoF.
-Therefore, the idea of using operational Jacobians in analyzing the teleoperation system’s robustness against manipulator singularities can also be generalized to teleoperation systems with different kinematic structures.
-Figure 5: Hardware setup of the Global–Local (G–L) teleoperation interface.
-(A), (B): Global teleoperation subsystem using a 3-DOF joint-link chain as the master device.
-(C), (D): Local teleoperation subsystem using a commercial haptic device as the master component.
+3 shows a schematic diagram of spatial decoupling.
+This decoupling architecture allows simultaneous activation of both components.
+The operator can manipulate the global component with one hand and the local component with the other.
+While seemingly adding complexity, this approach offers two key advantages.
+First, the dual-IMU design in the local component eliminates mechanical constraints around the operator’s hand, allowing integration of additional devices to control more sophisticated end-effectors.
+Second, replicating the first N3N-3 joints enables full control of the slave robot’s redundant DOFs, maximizing workspace utilization.
+Existing teleoperation systems often prioritize one of these advantages.
+Systems focusing on precise hand pose tracking [3, 2, 31] may have limited control over the full robot workspace.
+Conversely, systems emphasizing workspace utilization [18, 35, 1] may impose constraints on the operator’s hand, potentially hindering dexterity and the use of additional tools.
+Spatial decoupling combines both advantages, as demonstrated in the experiment section where a glove-shaped device is mounted on the local component to teleoperate a multi-fingered hand.
+Figure 4: Physical realization of temporal decoupling. (A), (B) Teleoperation with global component. (C), (D) Teleoperation with local component.
+Figure 5: Physical realization of spatial decoupling.
+(A) the operator moves the global and local components simultaneously.
+(B) The operator moves the local component only to manipulate the last three joints of the Flexiv robot arm.
+III Implementation Details
+Figure 6: The two-finger exoskeleton design in the spatial decomposition framework.
+The thumb of the Inspire hand follows the human operator’s thumb, and the index, middle, ring, and pinky fingers follow the human operator’s index.
+(B) shows how the two MTI-3 sensors are mounted on the operator’s arm for wrist rotation measurement.
+Figure 4 and 5 depicts the two teleoperation systems developed to implement the proposed temporal and spatial decoupling framework.
+The dual-arm teleoperation system, designed according to temporal decoupling, utilizes two Piper Agile X Limit robotic arms as the slave system.
+For gross manipulation, the global component employs Dynamixel servomotors, while the local component integrates two Touch X haptic devices for enhanced fine control.
+Each Piper arm, a 6-DOF lightweight manipulator, is equipped with a two-fingered gripper for position-controlled grasping.
+The master system’s global component comprises six Dynamixel servomotors: two XL430 motors for the base joints and four XL330-M288 motors for the remaining degrees of freedom.
+The Touch X devices, serving as the local component, are 6-DOF haptic interfaces that track the human operator’s finger tip position and orientation precisely.
+Similar to the single-arm system, all devices are managed and coordinated through a multi-process LabVIEW program.
 Figure 7: Snapshots of tasks requiring precise manipulation.
@@
 Examples (B), (C), (D), and (E) feature bimanual fine manipulation, where both slave arms need to maintain precise position or orientation for completion.
-VI Experimental Validations
-This section presents experimental results validating the performance of G-L interface leveraging the direct motion mapping (Section III) and singularity-avoiding motion mapping (Section IV).
-The evaluation focuses on three key aspects:
-  1. 1.
-    Precision Control: Using direct motion mapping, we assessed the system’s capability for high-accuracy telemanipulation through tasks with stringent precision requirements.
-  2. 2.
-    Workspace Coverage: Also employing direct motion mapping, we evaluated the ability to command extensive movements across the slave robot’s workspace, demonstrating effective utilization of its kinematic limits.
-  3. 3.
-    Orientation Dexterity and Singularity Handling: Utilizing the singularity-avoiding mapping, we demonstrated intuitive control during tasks involving significant orientation changes, specifically highlighting smooth trajectory execution during the traversal of wrist singularities.
-In addition to these demonstrations, a quantitative comparison was conducted to assess the benefits of the proposed integrated Global-Local (G-L) interface.
-Task completion times for both fine-grained manipulation and large-range motion scenarios were measured using the full G-L interface and compared against performance when using only the global or local component in isolation.
-This analysis serves to quantify the advantages derived from combining these complementary components.
-VI-A Implementation Details
-The global components are actuated with Dynamixel XL330-M288 servomotors arranged to a desired joint-link chain mimicking the Pieper robot arms.
-For direct mapping, the global component has full 6 DoFs, and for local mapping, it has only 3 DoFs.
-Fig.
-5 shows snapshots of G-L interface realization.
-The motion correspondence of the direct mapping and sigularity-avoiding realizations can be found in the supplementary video.
-The local component comprises two 3D System 6-DoF Touch devices for precise end-effector control in Cartesian space.
-The isolated global components in the comparison experiments has 6 DoFs.
-VI-B Precise Manipulation Demo
+The single-arm teleoperation system was developed based on spatial decoupling.
+Its slave system comprises a Flexiv Rizon robotic arm equipped with an Inspire dexterous hand.
+The global component of the master device utilizes a Dynamixel servomotor structure to control the robot arm’s primary joints, while the local component employs two MTI-3 inertial measurement units (IMUs) and a dexterous exoskeleton for fine manipulation.
+The Inspire hand, a 6-DOF, five-fingered anthropomorphic hand, is mounted to the Flexiv arm’s end-effector with a 90-degree offset to mitigate kinematic singularities.
+The master system’s global component consists of four serially-linked Dynamixel XL330-M288 servomotors, kinematically replicating the Flexiv arm’s proximal four joints.
+For the local component, two MTI-3 IMUs are attached to the operator’s forearm and palm, respectively, and a dexterous exoskeleton, integrated with an operational glove, captures index and thumb motions.
+All devices are interfaced and coordinated through a multi-process LabVIEW program.
+Upon first launch, the starting orientation of the two IMUs will be respectively recorded as the home orientation, denoted as $R^{i}_{1}$ and $R^{i}_{2}$.
+When the local component is activated and the two IMUs report current orientations $R_{1}$ and $R_{2}$, we take their rotational displacements ${R^{i}_{1}}^{-1}R_{1}$ and ${R^{i}_{2}}^{-1}R_{2}$.
+The difference between the forearm and hand IMU, which is $R_{s}=R_{1}^{-1}R^{i}_{1}{R^{i}_{2}}^{-1}R_{2}$, represents the combined rotation of the last three joints in the Flexiv arm.
+The last three joint positions are computed by finding the instrinsic X-Y-Z Euler angles that lead to $R_{s}$.
+An exoskeleton is designed for the single-arm teleoperation system to control the Inspire Hand.
+It is equipped with six AS5600 magnetic encoders and can be attached to the operator’s hand to record the thumb and index flexion of the human operator, as shown in Fig.
+6.
+The thumb of Inspire Hand will copy the motion of the operator’s thumb, whereas the index, middle, ring, and pinky fingers will imitate the motion of the operator’s index following the virtual finger concept [37, 38].
+The schematics and mapping methods of the exoskeleton will be detailed in the supplementary materials.
+IV Experimental Validations
+Figure 8: Comparison of completion time between global component only, local component only, and our proposed G-L interface.
+The G-L interface outperforms only global component in most tasks, except in the Nested Bucket Picking.
+As shown in Table II, despite taking less time to complete successful trials, only global component has a significantly lower success rate in Nested Bucket Picking.
+In large distance motion, G-L interface significantly outperforms local component only.
+The performance of the implemented temporal decoupling and spatial decoupling teleoperation systems, mentioned in section II, are evaluated through three distinct task categories.
+The first category comprises tasks with stringent accuracy requirements designed to assess the master system’s capacity for precise telemanipulation.
+The second category encompasses tasks necessitating extensive movement relative to the slave system’s workspace, demonstrating the system’s ability to effectively utilize the robot’s full range of motion.
+Finally, the third category involves tasks demanding the control of specialized dexterous end-effectors, specifically a five-fingered robotic hand in this study.
+IV-A Precise Manipulation
 Figure 9: Snapshots of tasks requiring large workspace.
@@
 (C) Bimanual handling of oranges from the front-left of the workspace to the back-right.
-(D) Bimanual transferring a test tube from a wash basin to a rack.
-(E) Pouring the liquid from a test tube into a beaker, and then place the test tube back on the rack.
-To assess the precision and dexterity afforded by the direct mapping teleoperation realization, we conducted a series of experiments involving intricate manipulation tasks.
+The supplementary video of this paper shows the whole process of these demonstrations.
+To assess the precision and dexterity afforded by the temporal decoupling teleoperation system (dual-arm setup), we conducted a series of experiments involving intricate manipulation tasks.
 These tasks, detailed below, were selected to represent challenges commonly encountered in real-world scenarios and to highlight the system’s capabilities in addressing them.
 Snapshots of these tasks are shown in Fig.
-6.
-Figure 8: Snapshots of daily object bimanual manipulation. (A) Opening a snack box. (B) Opening bottle cap. (C) Opening chip bag.
-Figure 9: Comparison of completion time between global component only, local component only, and our proposed G-L interface.
-The G-L interface outperforms only global component in most tasks, except in the Nested Bucket Picking.
-As shown in Table I, despite taking less time to complete successful trials, only global component has a significantly lower success rate in Nested Bucket Picking.
-In large distance motion, G-L interface significantly outperforms local component only.
+7.
   • Key Insertion: This task, known for its demanding precision requirements, involves inserting a key into a standard bronze lock.
@@
 This task underscores the system’s suitability for tasks requiring both precision and tool manipulation.
-VI-C Large Linear Range Manipulation Demo
-To assess the ability of the G-L interface to effectively utilize the entire workspace of the slave robot, we conducted two demonstrations involving large-scale movements spanning the workspace’s full range along the left-right, front-back, and vertical axes.
-These demonstrations aimed to confirm that the integration of precise manipulation capabilities within the G-L interface did not compromise its capacity for executing large-range manipulation tasks.
-  • Object Handling: This demonstration involved a coordinated bimanual task in which one slave robot retrieved an orange from the far left-front corner of its workspace and transferred it to the second slave robot.
-The second robot then handed off the orange to a human operator located at the far right-back corner of the workspace.
-This task exemplifies the system’s ability to facilitate long-range object transfer and cooperative manipulation, similar to demonstrations in prior teleoperation systems [29].
-  • Pick and Place from a Cabinet: This demonstration involved interacting with a tall cabinet positioned behind the slave robot.
-The task required retrieving specific objects from a cluttered arrangement within the cabinet and placing them in front of the robot, or conversely, placing objects from the workspace into the cabinet.
-This task highlights the system’s capacity for navigating and manipulating objects within complex, spatially extended environments.
-TABLE I: Success rates of isolated global component and G-L interface across different tasks. Tasks are selected based on predefined failure criteria described in Section VI-B.
-Task Global Component G-L Interface
-Nested Bucket Picking $3/7$ $6/7$
-Pencil Picking $/7$ 7 $/7$ 7
-Membrane Touching $5/7$ $/7$ 7
-Buckle Opening $6/7$ $/7$ 7
-Handle Lifting $/7$ 7 $/7$ 7
-Both tasks are completed successfully and their snapshots are shown in Fig.
-7.
-These tasks are good examples of how global and local components can cooperate to improve user experiences.
-The global component ensures a large range of motion while the local component enables the operator to make fine adjustments in clustered environments.
-VI-D Large Rotational Range Manipulation Demo
-To assess the ability of the singularity avoidance mapping on intuitive orientation control and robustness against wrist singularities, we conducted demonstrations involving considerable end-effector orientation change and needs to cross the wrist singularities.
-  • Test Tube Pick and Place: This demonstration involves picking up a flat-laying test tube from a basin, handle it over to another robot arm and place it on a rack.
-This task requires significant pitch angle change and passes through 156 and 456 singularities.
-  • Liquid Pouring: This demonstration involves picking up a test tube from the rack, pour the liquid into a beaker and place the test tube rack. This task requires significant yaw angle change.
-Both tasks are completed efficiently and their snapshots are shown in Fig.
-7D and 7E.
-These tasks are good examples of how singularity avoidance motion mapping yields intuitive orientation control and robustness against wrist singularities.
-VI-E Daily Object Manipulation
-We additionally show the G-L interface can also complement bimanual teleoperation tasks involving daily objects as in Fig.
-8.
-These tasks include picking up a plate, where one slave robot must push down one edge to lift up the other, thereby allow another slave robot to pick up the entire plate.
-Tasks like opening a chip bag, snack box, and water bottle are also included.
-VI-F Fine-Grained Task Completion Time Comparisons
-To quantitatively assess the performance improvements enabled by the G–L hierarchical teleoperation system, we conducted a comparative study involving three control conditions: (1) using the global component only, (2) using the local component only, and (3) using both the global and local components.
-Seven participants were recruited to perform a series of representative tasks, each completed twice under every condition.
-Regrasping and retrying were allowed unless the task was classified as a failure based on predefined criteria (detailed below).
-Failed trials were excluded from the computation of average completion time.
-Prior to the experiment, all participants received a brief tutorial on system operation and were allotted one hour of practice to familiarize themselves with the interface and task procedures.
-For each trial, the master and slave systems were reset to identical initial configurations, and the task completion time was recorded.
+To quantitatively evaluate the performance improvements offered by the temporal decoupling teleoperation system, we conducted a comparative study with isolated global and local component.
+Seven participants were recruited to complete a series of tasks under two conditions: (1) using the global component (equivalent to the GELLO system) or local component only, and (2) using both the global and local components.
+Each participant performed each task twice, once under each condition.
+Regrasping and retrying were permitted unless the task was deemed failed according to the predefined criteria detailed below.
+In the event of failure, the respective trial was disregarded when computing the average completion time.
+Prior to the experiment, participants received a brief tutorial on operating the teleoperation system and were given one hour of practice time to familiarize themselves with the system and task procedures.
+For each trial, the master and slave systems were initialized to identical starting configurations, and the completion time was recorded.
 Snapshots of the following tasks designed for assessment can be found in the supplementary materials.
@@
 The task was considered failed if the emergency stop is triggered.
-  • Handle Lifting: Participants rotate a laid-down box handle to an upright position and lift the box. Failure conditions are identical to the Buckle Opening task.
-  • Dots Aiming: Abstracting common insertion tasks (e.g., SIM ejector, shafts), participants use the slave gripper to prick three target dots (1 mm and 3 mm radius) with a needle.
-The target sheet’s location is hidden during trials but available for practice.
-No failure criterion; continuous attempts are allowed until success.
-  • Large Distance Motion: This task compares the G–L interface with a local-only teleoperation setup to evaluate efficiency in executing large-range movements.
-When repeatedly performing such motions, the G–L interface demonstrates a clear advantage in operational efficiency.
-Specifically, the G–L interface achieves a significantly shorter task completion time, averaging only $3.30\pm 0.53$ seconds, compared to $15.1\pm 1.3$ seconds using the local component alone.
+  • Handle Lifting: Given a box whose handle is initially laid down, participants are required to rotate the handle to an upright pose and lift the box using the handle.
+The failure condition is the same as the Buckle Opening task.
+  • Dots Aiming: This task is an abstraction of many accurate insertion tasks like SIM card needles, shafts, and TRS connectors.
+Three dots with radius 1 mm, 2 mm, and 3 mm are printed on a target sheet.
+The slave system pinches a needle with its gripper.
+Volunteers are asked to prick each dot with this needle using the master system.
+Participants cannot access the test time target sheet position, but they are allowed to practice with the same target sheet before the test starts.
+This task is not given a failure criteria so participants can make continuous attempts until they are successful.
+Figure 10: Snapshots of daily object manipulation. (A) Opening a snack box. (B) Dual arm plate grasping. (C) Opening bottle cap. (D) Opening chip bag.
 Fig.
@@
 The isolated global component system demonstrated marginally faster completion in nested bucket picking.
-However, it has a considerably lower success rate comparing to the G-L interface, as elaborated in Table I.
-VII Conclusion, Limitations, and Future Work
-This paper presents the Global–Local (G–L) teleoperation interface, which decouples large-scale positioning and fine manipulation through a hierarchical dual-mode control framework.
-We implemented both direct and singularity-avoiding motion mapping strategies, and introduced the concept of an operational Jacobian to analyze the smoothness and stability of joint behavior during local control.
-Experimental results across a range of tasks—including precision placement, large workspace traversal, and bimanual manipulation—demonstrate that the G–L interface improves efficiency, task success rate, and user experience compared to conventional global-only or local-only systems.
-Nonetheless, several limitations remain.
-Manual mode switching introduces operational friction and may increase cognitive load.
-The current singularity-avoiding mapping strategy is task-driven and heuristic, lacking general optimality or theoretical uniqueness.
-Additionally, the system does not yet support force–torque feedback, and has not been validated under remote or delayed network conditions.
-Furthermore, the G–L interface’s modular architecture supports flexible combinations of global and local master devices, making it broadly adaptable to various application scenarios beyond the current implementation.
-Future work will focus on integrating automatic or learning-based switching mechanisms, increasing the system’s degrees of freedom to improve singularity robustness, and incorporating visual and force–torque feedback for closed-loop, compliant control.
-We also plan to extend the operational Jacobian framework to accommodate non-Pieper structures and redundant manipulators, enabling broader applicability and more adaptive mapping strategies.
-These developments aim to further enhance the robustness, responsiveness, and scalability of the G–L teleoperation interface for real-world deployment in industrial, surgical, and exploratory environments.
+However, it has a considerably lower success rate comparing to the G-L interface, as elaborated in Table II.
+Global Component G-L Interface
+Nested Bucket Picking $3/7$ $6/7$
+Pencil Picking $7/7$ $7/7$
+Membrane Touching $5/7$ $7/7$
+Buckle Opening $6/7$ $7/7$
+Handle Lifting $7/7$ $7/7$
+TABLE II: Success rate of isolated global component and G-L interface on each task.
+Only tasks with pre-defined failure conditions in section IV-A are included.
+The G-L interface shows a higher success rate
+IV-B Large Range Manipulation
+Figure 11: Snapshots of tasks with dexterous manipulation.
+(A) The operator needs to fetch a box of drink from a distance away, take a sip and then put the drink box back.
+(B) Reorientating a water bottle from the desk.
+To assess the ability of the G-L interface to effectively utilize the entire workspace of the slave robot, we conducted two demonstrations involving large-scale movements spanning the workspace’s full range along the left-right, front-back, and vertical axes.
+These demonstrations aimed to confirm that the integration of precise manipulation capabilities within the G-L interface did not compromise its capacity for executing large-range manipulation tasks.
+  • Object Handling: This demonstration involved a coordinated bimanual task in which one slave robot retrieved an orange from the far left-front corner of its workspace and transferred it to the second slave robot.
+The second robot then handed off the orange to a human operator located at the far right-back corner of the workspace.
+This task exemplifies the system’s ability to facilitate long-range object transfer and cooperative manipulation, similar to demonstrations in prior teleoperation systems [1].
+  • Pick and Place from a Cabinet: This demonstration involved interacting with a tall cabinet positioned behind the slave robot.
+The task required retrieving specific objects from a cluttered arrangement within the cabinet and placing them in front of the robot, or conversely, placing objects from the workspace into the cabinet.
+This task highlights the system’s capacity for navigating and manipulating objects within complex, spatially extended environments.
+Both tasks are completed successfully and their snapshots are shown in Fig.
+9.
+These tasks are good examples of how global and local components can cooperate to improve user experiences.
+The global component ensures a large range of motion while the local component enables the operator to make fine adjustments in clustered environments.
+IV-C Daily Object Manipulation
+Tasks demonstrated in sections IV-A and IV-B are mainly achieved via single-arm manipulation.
+In this section, we show that the temporal decoupling teleoperation system can also complement bimanual teleoperation tasks involving daily objects.
+These tasks include picking up a plate, where one slave robot must push down one edge to lift up the other, thereby allow another slave robot to pick up the entire plate.
+Tasks like opening a chip bag, snack box, and water bottle are also included.
+IV-D Dexterous Manipulation
+This section details the master system built from spatial decoupling that is capable of controlling a complex end-effector.
+We use two tasks to demonstrate its dexterity: Feeding the human operator a drink and reorientating a water bottle.
+Snapshots of these tasks are shown in Fig.
+11.
+V Discussions and Limitations
+Despite its advantages, the proposed G-L interface presents certain limitations, primarily stemming from its design and hardware implementation.
+These limitations, outlined below, warrant consideration for future development and refinement.
+  • Additional Hardware Costs: Unlike conventional teleoperation systems that typically employ a single master device per slave robot, the G-L interface often requires two sets of master devices.
+While this configuration contributes to enhanced performance, it also leads to increased hardware costs.
+  • Bimanual Control for Unimanual Tasks: The temporal decoupling method inherent to the G-L interface allows for straightforward bimanual teleoperation by preventing simultaneous activation of the global and local components.
+However, in the spatial decoupling method, where both components are engaged concurrently, the operator must utilize both hands to control a single slave robot, potentially increasing operational complexity.
+  • Challenges in Orientation Control: Empirical observations revealed that fine-tuning the orientation of the slave robot’s end-effector proved more challenging than adjusting its position.
+Within the temporal decoupling framework, the local component’s reliance on scaled motion tracking from the touch-X device limits its efficacy to small orientation adjustments.
+Larger adjustments can become unintuitive due to the scaled mapping between master and slave rotations.
+Conversely, while the dual-IMU design in the spatial decoupling framework permits larger orientation adjustments, it introduces sensitivity to unintended skin movement as the operator twists their arm.
+These movements can increase discrepancies between the measured wrist rotation and its ground truth, leading to counterintuitive behavior in the slave end-effector’s orientation.
+This issue could potentially be mitigated by incorporating a more robust and precise wrist rotation sensing mechanism.
+Interestingly, a notable learning curve was observed among participants.
+Initial trials were often characterized by cautious and slow movements, with proficiency and speed increasing with practice.
+To account for this learning effect and ensure a fair comparison, each participant was provided with one hour of dedicated practice time prior to data collection.
+VI Conclusion
+This paper introduced the G-L interface for on-demand teleoperation, a novel approach that decouples the master system into global and local components responsible for large-range movements and fine manipulations, respectively.
+This design enables the master system to effectively make precise adjustments while retaining the capacity for utilizing the entire workspace of the slave robot.
+The modular nature of the G-L interface allows for flexible implementation tailored to specific task requirements.
+This work presented two realizations of the G-L interface: temporal decoupling and spatial decoupling.
+Through demonstrations and experimental evaluation, we have shown that temporal decoupling facilitates tasks demanding high accuracy while maintaining full workspace utilization.
+Conversely, spatial decoupling enables the teleoperation of complex end-effectors with enhanced dexterity.
+These findings highlight the versatility and potential of the G-L interface to advance the capabilities of teleoperation systems across a wide range of applications.
+Appendix A: Exoskeleton Design
+Figure 12: Design details of the exoskeleton master that controls the Inspire dexterous hand.
+For better fit with the human operator’s hand, we added a $60^{\circ}$ tilting angle to the thumb of the exoskeleton.
+Fingers on the exoskeleton have two joints, each equipped with an AS5600 magnetic encoder.
+We designed a circuit to send encoder readings to a multiplexer via I2C protocol.
+The multiplexer is then connected to an ESP32 microprocessor, which allows communication with a computer.
+Fig.
+12 further elaborates the exoskeleton’s structural design and the communication circuit.
+Figure 13: Mapping between the exoskeleton and Inspire dexterous hand’s state. Subfigures (A), (B) and (C) are duplicated from the manual of the Inspire hand.
+The Inspire dexterous hand is cable-driven with 6 DoFs in total, two on the thumb and one for each of the remaining fingers.
+Its position state is represented by the bending and rotational angles between the fingertip and the metacarpal plane.
+Therefore, while mapping the exoskeleton’s position state to the Inspire hand’s position state, we also compute the angle between the fingertip and metacarpal plane of the exoskeleton, as shown in Fig.
+13.
+Appendix B: Further Details of Tasks in Fig. 8
+This part includes snapshots of tasks in Fig.
+8 of the paper for better understanding of the testing criteria.
+Tasks that have already appeared in Fig.
+7 and the dots aiming tasks are not included here.
+(a) Nested Bucket Picking: Taking out the smaller bucket from the larger bucket.
+Reattempts commonly happens when both buckets are taken up, or the larger bucket is tipped over when lifting up the smaller bucket.
+Reattempts usually cause an uncontrollable motion of the larger bucket, and the task is considered failed if the large bucket fall off the platform.
+(b) Pencil Picking: If the gripper hits the desk too hard, emergency stop will be triggered and the task is considered failed.
+(c) Membrane Touching: The middle subfigure is a succeed case where the needle is touching the scotch tape. The right subfigure is a failure case where the needle penetrates through the scotch tape.
+(d) Buckle Opening: If the gripper pushes the box too hard in a wrong direction, emergency stop will be triggered and the task is considered failed.
+(e) Handle Lifting: the slave robot must rotate the and and lift up the box to be considered success.
+Figure 14: Snapshots and further explanations of tasks for quantitative studies in Fig. 8.